### v4

# Fine-Grained Image-Text Alignment in Medical Imaging Enables Cyclic Image-Report Generation

Abstract
Fine-grained vision-language models (VLM) have been widely used for inter-modality local alignment between the predefined fixed patches and textual words. However, in the medical analysis, lesions exhibit varying sizes and positions, and the fixed patches may cause incomplete representations of lesions. Moreover, these methods provide explainability by using heatmaps to show the general potential image areas associated with texts rather than specific regions, making their explanations not explicit and specific enough. To address these issues, we propose a novel Adaptive patch-word Matching (AdaMatch) model to correlate chest X-ray (CXR) image regions with words in medical reports and apply it to CXR-report generation to provide explainability for the generation process. AdaMatch exploits the fine-grained relation between adaptive patches and words to provide explanations of specific image regions with corresponding words. To capture the abnormal regions of varying sizes and positions, we introduce the Adaptive Patch extraction (AdaPatch) module to acquire the adaptive patches for these regions adaptively. In order to provide explicit explainability for CXR-report generation task, we propose an AdaMatch-based bidirectional large language model for Cyclic CXR-report generation (AdaMatch-Cyclic). It employs the AdaMatch to obtain the keywords for CXR images and ‘keypatches’ for medical reports as hints to guide CXR-report generation. Extensive experiments on two publicly available CXR datasets prove the effectiveness of our method and its superior performance to existing methods.
1 Introduction
Inter-modality alignment, such as vision and language, has been an important task with growing interests in the field of computer vision, especially with the recent advancement in representation learning [34]. Technologies like contrastive learning and self-supervised learning have dramatically improved state-of-the-art alignment performance. Recent vision-language models (VLMs) demonstrate two approaches: global contrastive alignment, which integrates images and texts at a global level [34, 19, 17, 44, 46], and local alignment, focusing on detailed connections between visual objects and textual words [6, 25, 24, 52, 21, 50], as illustrated in Fig. 1.
Figure 1: Current vision-language models (VLM) achieve (a) global alignment and (b) local alignment by matching overall visual with textual features, and aligning patch with word features, respectively. (c) To exploit the relation between textual words and abnormal patches with different sizes, the proposed AdaMatch obtains adaptive patch features and aligns them with word features.
Current VLMs with local alignment either adopt the pre-trained object detector to extract region-of-interest (ROI) features from images and match the corresponding object features with textual words [6, 25, 24, 52]; or align the visual token from each patch and the textual token into the same embedding space [21, 50, 18, 41]. The former highly relies on the quality of the object detector and its predefined classes, which is less generalizable to new domains. The latter family of methods learns the alignment in a more automatic and data-driven manner. However, most of these methods depend on pre-defined patch size and position (e.g., grids) across images. In most challenging cases, such as the analysis of medical image analysis, lesions can exhibit a wide range of shapes, sizes, and positions. A fixed definition of image patches can lead to incomplete or ambiguous representations of the key imaging abnormalities. Hence, it is highly needed to adaptively exploit the fine-grained relationship between image embeddings derived from a more flexible patching scheme and textual embeddings.
Another challenge in the current local medical VLMs lies in their explainability: it is generally difficult to delineate the image-text relationship learned by the model, especially for the current medical VLMs. Current solutions to provide such explanations in medical VLMs leverage the attention maps from the intermediate layer to visualize the location of the abnormalities [32, 16, 45]. Other methods  [40, 5, 28] utilize network gradients such as Grad-CAM [38] to generate the heatmaps according to lesion types based on ground-truth reports. However, both maps can only show the general potential areas associated with the corresponding text data rather than clearly pinpointing a specific region. In addition, gradient-based methods need ground-truth reports, prohibiting them from functioning correctly beyond training data. It is thus highly necessary to develop a mechanism that could provide explicit and specific explanations of the input image or text during inference time.
Figure 2: The overview of the proposed methods. (a) Adaptive patch-word Matching (AdaMatch) model. (b) AdaMatch-based bidirectional large language model (LLM) for cyclic CXR-report generation (AdaMatch-Cyclic).
To address these two challenges above, we propose a novel Adaptive patch-word Matching (AdaMatch) model to match fine-grained image regions of various sizes and positions with textual data. AdaMatch introduces an image encoder with multiple Adaptive Patch extraction (AdaPatch) modules to adaptively acquire the patches associated with certain text tokens. It then performs patch-word alignment based on contrastive learning. AdaMatch is specifically developed in the context of aligning radiology images (chest X-ray, CXR) and their corresponding radiology reports with the capability of achieving cyclic (CXR-to-report and report-to-CXR) generation based on the learned alignment. Our premise is that such a cyclic generation task would serve as the best use case and evaluation criterion for the desired fine-grained alignment. Also, the fine-grained cyclic generation between CXR and report will provide a natural explainability for how the model aligns the two modalities: for any given text token, we can visualize its matching imaging manifestation; and for any image region within a CXR image, we can tell the type of lesion or anatomical region it belongs to.
To implement the cyclic CXR-report generation, we propose an AdaMatch-based bidirectional model (AdaMatch-Cyclic). AdaMatch-Cyclic employs AdaMatch to identify the keywords for CXR images and ‘keypatches’ for medical reports to guide the generation tasks. Since the potential keywords for CXR images cover a wide range and ground-truth reports cannot be used during inference, we predefine a textual codebook with the most common entities from medical reports as prior knowledge during fine-grained alignment. With the textual codebook, AdaMatch aligns it with the adaptive patches to obtain matched keywords as guidance for report generation. Next, a VQ-GAN model encodes the CXR image into image tokens, and a Large Language Model (LLM) takes image tokens, the matched keywords, and the instructions as input to generate the medical reports. Similarly, we also build a visual codebook with the most commonly seen patches as ‘keypatches’, and use AdaMatch to obtain the matched keypatches from given text reports as hints for CXR generation. With medical reports, matched keypatches, and instructions, LLM outputs the image tokens, and then the VQ-GAN model decodes them into the generated CXR image. Our contributions are summarized as follows:
  • To exploit the fine-grained relation between the CXR image patches with words of medical reports, we propose an Adaptive patch-word Matching (AdaMatch) model to obtain adaptive patches for abnormal regions and perform alignment between them and texts in medical reports.
  • We devise an AdaMatch-based bidirectional LLM for Cyclic CXR-report generation (AdaMatch-Cyclic) to facilitate the generation between CXR and reports. Moreover, we build the textual and visual codebook to utilize AdaMatch to extract useful keywords and keypatches for the report and CXR generation, respectively.
  • Experiments on two publicly available chest X-ray datasets demonstrate the effectiveness of our method and its superior performance over state-of-art methods.
2 Related Works
2.1 Fine-Grained Vision-Language Models
Recently, several fine-grained vision-language models (VLM) [6, 25, 24, 52, 21, 50, 18, 41] achieve the local alignment by exploiting the fine-grained relation between visual objects and textual words. Some methods [6, 25, 24, 52] employ the pre-trained object detector to obtain region-of-interest (ROI) features from images and align the object features with textual features, and others [21, 50, 41, 18] aim to locally align the fixed patches with the textual words. For instance, Li et al. [24] utilizes Faster-RCNN [36] to detect visual objects and match ROI embeddings with textual embeddings. Moreover, FILIP [50] achieves cross-modal interaction by maximizing the similarity between the visual tokens of fixed patches and textual tokens. The former requires a precise object detector, and the latter primarily learns the relation between fixed patches inside the predefined grid and words. Nevertheless, when lung lesions differ in size, shape, and position, these methods may split them into separate patches, leading to incomplete semantic information of patches. Thus, we devise an Adaptive patch-word Matching (AdaMatch) model to adaptively exploit the fine-grained relation between flexible patches and textual words.
2.2 CXR-Report Generation
Medical VLMs are widely used in downstream tasks for chest radiographs, including CXR-to-report generation [7, 8, 48, 42, 39, 47] and report-to-CXR generation [37, 3, 2, 22, 23] tasks. CXR-to-report generation has been investigated to reduce the burden of radiologists and promote clinical automation. For instance, Chen et al. [8] introduces a cross-modal memory network with shared memory to align images with texts to promote report generation performance. In report-to-CXR generation task, previous methods generate high-quality and annotated CXR images from medical reports, to provide more training data and avoid privacy laws and considerations. Some diffusion-based methods [2, 22] take the impression section of reports as input, ignoring the finding section that includes detailed descriptions. To consider more details in reports, several transformer-based methods [22, 23] take both finding and impression sections as input to synthesize CXR. In this paper, we use the cyclic generation (i.e. CXR-to-report and report-to-CXR generation) as use case and evaluation criteria for our fine-grained alignment method (AdaMatch). We also design an AdaMatch-Cyclic to employ AdaMatch to improve the explainability of cyclic generation.
3 Methods
In Fig. 2, we propose an adaptive patch-word matching (AdaMatch) model to associate the CXR image regions with words in medical reports, and apply it to CXR-report generation to provide explainability. Given an input CXR image $x^{I}$, the image encoder with several Adaptive Patch extraction modules (AdaPatch) predicts the location and scale of multiple adaptive patches and processes the adaptive patch embeddings $z^{I}$. With $z^{I}$ and text embeddings $z^{T}$ extracted by text encoder for the medical report $x^{T}$, we compute the similarities $s^{I}$, $s^{T}$ between adaptive patches and text tokens to calculate the contrastive loss $L_{I2T},L_{T2I}$ to optimize AdaMatch model. For CXR-to-report generation, we utilize the frozen AdaMatch model to match a predefined textual codebook with the input CXR image $x^{I}$ to obtain matched keywords $k^{T}$ and feed a LLM with $k^{T}$, the instruction, and image tokens encoded by VQ-GAN from $x^{I}$ to generate a medical report $x^{T^{\prime}}$. Then, the AdaMatch model is used to match a predefined visual codebook with the generated report $x^{T^{\prime}}$ to acquire the matched keypatches $k^{I}$. With instruction, $x^{T^{\prime}}$ and $k^{I}$, LLM outputs image tokens of the generated CXR image $x^{I^{\prime}}$, which is optimized by $L_{text}$ and $L_{img}$.
3.1 Adaptive Patch-Word Matching (AdaMatch)
Current VLMs [6, 25, 24, 52, 21, 50] achieve fine-grained alignment between visual objects and textual words, but they may split lung lesions into separate fixed patches due to various shapes, sizes, and positions of lung lesions. Thus, we propose an Adaptive patch-word Matching (AdaMatch) to automatically locate the important regions, extract adaptive patches for these regions, and align them with corresponding textual words. As shown in Fig. 2 (a), AdaMatch includes an image encoder to extract adaptive patch features and a text encoder to encode textual features.
Image Encoder. To obtain the adaptive patches, the image encoder consists of four stages with feature maps of decreasing scales. Specifically, in the first stage, we first adopt the patch embedding module to split the CXR image $x^{I}\in\mathbb{R}^{H\times W\times C}$ into $N$ patches with fixed size $s\times s$ and project them to patch embeddings $z^{(i)}(1\leq i\leq N)$ through a fully connected layer $g$, $z^{(i)}=g([a^{(i,1)};\dots;a^{(i,s\times s)}])$, where $a^{(i,j)}$ indicates the image features for the pixel located at $q^{(i,j)}$ and []delimited-[][$\cdot$][ ⋅ ] denotes the concatenation operation. Afterward, we pass the patch embeddings $z^{(i)}$ with the position embeddings $e_{pos}$ into a transformer encoder [43] $R$ to obtain the outputs, $z^{(i)}=R(z^{(i)}+e_{pos})$, where $R$ includes an attention layer and a feed-forward layer. In order to locate the potential lung lesions, we devise an Adaptive Patch Extraction module (AdaPatch) for the rest of stages. In AdaPatch, a fully connected layer is fed with patch embeddings $z^{(i)}$, and then the outputs are passed through four separate fully connected layers $f_{1},f_{2},f_{3},f_{4}$ to predict the offset $(\delta_{x}^{(i)},\delta_{y}^{(i)})$ and patch size $(s_{w}^{(i)},s_{h}^{(i)})$ for the adaptive patches, where the offset indicates the shift to the center of original fixed patches $(c^{(i)}_{x},c^{(i)}_{y})$,
$\delta_{x}^{(i)}=\small{Tanh}(f_{1}(z^{(i)})),\delta_{y}^{(i)}=\small{Tanh}(f_{2}(z^{(i)})),$ (1)
$s_{w}^{(i)}=\small{ReLU}(\small{Tanh}(f_{3}(z^{(i)})),s_{h}^{(i)}=\small{ReLU}(\small{Tanh}(f_{4}(z^{(i)})).$ (2)
With the offset and patch size, we compute the position of left-top $(a^{(i)}_{x},a^{(i)}_{y})$ and right-bottom corners $(b^{(i)}_{x},b^{(i)}_{y})$ for each adaptive patch,
$a^{(i)}_{x}=c^{(i)}_{x}+\delta_{x}^{(i)}-\frac{s_{w}^{(i)}}{2},a^{(i)}_{y}=c^{(i)}_{y}+\delta_{y}^{(i)}-\frac{s_{h}^{(i)}}{2},$ (3)
$b^{(i)}_{x}=c^{(i)}_{x}+\delta_{x}^{(i)}+\frac{s_{w}^{(i)}}{2},b^{(i)}_{y}=c^{(i)}_{y}+\delta_{y}^{(i)}+\frac{s_{h}^{(i)}}{2},$ (4)
and uniformly sample $m\times m$ feature points inside the patches. Since the coordinates may be fractional, bilinear interpolation is used to obtain the sampled feature points $\left\{\hat{p}^{(j)}\right\}_{1\leq j\leq m\times m}$. The embeddings of all sampled points are flattened and fed into a fully connected layer $f_{5}$ to obtain patch embeddings, $z^{(i)}=f_{5}([\hat{p}^{(i,1)};\dots;\hat{p}^{(i,m\times m)}])$. Finally, we pass $z^{(i)}$ with position embeddings $e_{pos}$ into a transformer encoder $R$. The final adaptive patch embeddings $z^{I}\in\mathbb{R}^{N\times d}$ are the ensemble of $z^{(i)}(1\leq i\leq N)$, $z^{I}=E^{I}(x^{I})=\left\{z^{I}_{1},\cdots,z^{I}_{N}\right\},$ where $E^{I}$, $x^{I}$, and $d$ denote the image encoder, the input CXR image, and the dimension of adaptive patch embeddings, respectively.
Text Encoder. We adopt the BioClinicalBERT [1] as the text encoder $E^{T}$ to encode the medical report $x^{T}$ into text embeddings $z^{T}$. To make $z^{T}$ with the same dimension as $z^{I}$, we feed a fully connected layer with $z^{T}$ to reduce its dimension, $z^{T}=E^{T}(x^{T})=\left\{z^{T}_{1},\cdots,z^{T}_{K}\right\}(z^{T}\in\mathbb{R}^{K\times d})$, where $K$ denotes the number of text tokens.
Patch-Word Alignment. In order to exploit the relation between the adaptive patches and textual tokens, we perform fine-grained contrastive representation learning to achieve patch-word alignment. Concretely, for the $i$-th CXR image $x^{I}_{i}$ and $j$-th medical report $x^{T}_{j}$, we first compute the similarities between all the adaptive patch embeddings $z^{I}_{n}(1\leq n\leq N)$ and all the text embeddings $z^{T}_{k}(1\leq k\leq K)$, and use the largest similarity $\max_{(1\leq k\leq K)}{{(z^{I}_{n})}^{\top}z^{T}_{k}}$ as the patch-word maximum similarity for $n$-th adaptive patch embedding. Then, the patch-word maximum similarities for all the adaptive patch embeddings are averaged as the similarity $s^{I}_{i,j}$ of the $i$-th CXR image to the $j$-th medical report,
$s^{I}_{i,j}=\frac{1}{N}\displaystyle\sum_{n=1}^{N}(z^{I}_{n})^{\top}z^{T}_{m_{n}^{I}},$ (5)
where $m_{n}^{I}=\arg\max_{(1\leq k\leq K)}{{(z^{I}_{n})}^{\top}z^{T}_{k}}$. Similarly, the similarity of the $j$-th medical report to the $i$-th CXR image is defined as,
$s^{T}_{i,j}=\frac{1}{K}\displaystyle\sum_{k=1}^{K}(z^{I}_{m_{k}^{T}})^{\top}z^{T}_{k},$ (6)
where $m_{k}^{T}=\arg\max_{(1\leq n\leq N)}{{(z^{I}_{n})}^{\top}z^{T}_{k}}$. We exclude the padded textual tokens when computing the similarity.
With the cross-modal similarities $s^{I}_{i,j}$ and $s^{T}_{i,j}$ for the $i$-th CXR image and $j$-th medical report, we compute the adaptive patch-word contrastive loss $L_{i}^{I}$ for $i$-th CXR image $x_{i}^{I}$,
$L_{i}^{I2T}(x_{i}^{I},\left\{x_{j}^{T}\right\}_{j=1}^{b})=-\frac{1}{b}\log\frac{exp(s^{I}_{i,i}/\tau)}{\textstyle\sum_{j}{exp(s_{i,j}^{I}/\tau)}},$ (7)
where $b$ denotes the batch size, $\tau$ represents temperature hyperparameter and $\left\{x_{i}^{I},x_{i}^{T}\right\}$ indicates the positive CXR-report pair. Similarly, the adaptive word-patch contrastive loss $L_{i}^{T}$ for $i$-th medical report $x_{i}^{T}$ is formulated as,
$L_{i}^{T2I}(x_{i}^{T},\left\{x_{j}^{I}\right\}_{j=1}^{b})=-\frac{1}{b}\log\frac{exp(s^{T}_{i,i}/\tau)}{\textstyle\sum_{j}{exp(s_{j,i}^{T}/\tau)}}.$ (8)
The final contrastive loss for a mini-batch is calculated by,
$L=\frac{1}{2}\sum_{i=1}^{b}{(L_{i}^{I2T}+L_{i}^{T2I})}.$ (9)
With the final contrastive loss, the image and text encoders are optimized to automatically locate the important patches with varied sizes and exploit the fine-grained relation between the patches and words.
3.2 AdaMatch-based LLM for Cyclic CXR-Report Generation (AdaMatch-Cyclic)
To provide explicit explainability for CXR-report generation task, we propose an AdaMatch-based bidirectional LLM for the cyclic CXR-report generation (AdaMatch-Cyclic) by locating the potential lesions in CXR images and visualizing the appearance of description in reports to guide the generation process, as depicted in Fig. 2 (b).
Figure 3: Instruction data example for CXR-to-report generation.
3.2.1 CXR-to-Report Generation
In CXR-to-report generation, we employ the AdaMatch to match a predefined textual codebook with the CXR image to obtain the matched keywords as hints to guide LLM to generate reports.
Building textual codebook. Specifically, to establish a textual codebook, we use a pre-trained BioEN [35] to extract the related entities from the medical reports in the training set, where the related entities are divided into four entity groups, i.e., biological structure, detailed description, disease disorder, and sign symptom. Next, we compute the frequency of each entity and pick top $\kappa_{0}$ entities for each entity group as keywords in the textual codebook.
Keywords Extraction. With the textual codebook, we employ the frozen AdaMatch model to match the keywords from the textual codebook with the adaptive patches of CXR images, and obtain a patch-word similarity matrix $s^{P}\in\mathbb{R}^{M\times N}$ between the keyword tokens and adaptive patch embeddings, where $M$ and $N$ denote the number of keyword tokens and adaptive patches. To extract the most matched keywords for each adaptive patch, we rank the patch-word similarity along the dimension of keyword tokens, obtain the top $\kappa_{1}$ patch-word similarities for each adaptive patch, and extract the corresponding keywords $k^{T}$. The matched keywords can explain potential lesions in each adaptive patch, to assist LLM in generating medical reports.
Figure 4: Instruction data example for report-to-CXR generation.
Instruction tuning LLM. After keywords extraction, we use a frozen VQ-GAN [13] encoder $E$ to encode the input CXR image $x^{I}$ to the quantized image latent vectors as image tokens $E(x^{I})$. For CXR-to-report generation, the dolly-v2-3b [9] model is adopted as the bidirectional LLM, which is pre-trained on the instruction-following dataset. When fine-tuning LLM, we convert CXR-to-report generation dataset in instruction-following format. An example of instruction data for CXR-to-report generation is shown in Fig. 3, including the instruction, input, keywords, and response parts, where input is the image tokens of the input CXR image, keywords are extracted by AdaMatch, and response is the ground-truth medical report. During training, the response part is hidden and the LLM learns to generate it in autoregressive manner.
By adopting AdaMatch-Cyclic for CXR-to-report generation, we can locate the potential lesion in CXR images, to provide more explainability for the generation process.
3.2.2 Report-to-CXR Generation
In report-to-CXR generation, we leverage the AdaMatch to match a predefined visual codebook with the generated reports $x^{T^{\prime}}$ to acquire matched keypatches as guidance for LLM to synthesize CXR images, where ‘keypatches’ are important image patches related to generated reports.
Building visual codebook. Concretely, we first construct a visual codebook with the most common adaptive patches as keypatches. To collect the most common adaptive patches for CXR images in the training set, the AdaMatch is utilized to match the adaptive patches of CXR images with textual tokens of medical reports to obtain top $\kappa_{2}$ CXR-report pairs with the highest report-to-CXR similarities $s^{T}$. For each CXR-report pair, we compute the word-patch maximum similarity $\max_{(1\leq n\leq N)}{{(z^{I}_{n})}^{\top}z^{T}_{k}}$ for each textual token, rank the word-patch maximum similarities, and extract the adaptive patches for top $\kappa_{3}$ similarities as keypatches in the visual codebook. Each keypatch includes its adaptive patches and the corresponding features.
Keypatches Extraction. With the visual codebook, the frozen AdaMatch matches the features of keypatches in the visual codebook with textual tokens of the generated report to acquire the word-patch similarity matrix $s^{W}\in\mathbb{R}^{(\kappa_{2}\times\kappa_{3})\times K}$, where $K$ denotes the number of textual tokens for the generated report. To obtain keypatches related to the generated report, we rank the word-patch similarity along the dimension of keypatches, obtain the top $\kappa_{4}$ word-patch similarity for each textual token, and extract the features of corresponding keypatches $k^{I}$.
Instruction tuning LLM. After keypatches extraction, we employ a frozen VQ-GAN encoder to convert the matched keypatches $k^{I}$ into image tokens $E(k^{I})$, and feed the pre-trained LLM with the instruction, the generated report, and image tokens of keypatches in the instruction-following format, as shown in Fig. 4. Then, LLM predicts the image tokens and the VQ-GAN decoder decodes the image tokens into the corresponding generated CXR image $x^{I^{\prime}}$.
AdaMatch-Cyclic allows us to interpret generated reports with matched keypatches, thereby providing explainability to the generation procedure.
3.2.3 Overall Objective
To optimize AdaMatch-Cyclic, we apply the standard language modeling objective for both CXR-to-report and report-to-CXR generation. For CXR-to-report generation, LLM is fed with the instruction $I^{t}$, image tokens of input CXR image $E(x^{I})$, and matched keywords $k^{T}$ in instruction-following format, and LLM aims to generate the ground-truth medical report of response part autoregressively. We compute the $P(u_{k})$ conditional probability of the $k$-th token, $P(u_{k})=\mathcal{LLM}(I^{t},E(x^{I}),k^{T})$, where $k$ is the token index after the response key (### Response:). The report generation loss $L_{text}$ for the response area is calculated by,
$L_{text}=\sum_{i=k}^{n}-\log P(u_{i}|u_{1},u_{2},\cdots,u_{i-1}),$ (10)
where $[u_{1},u_{2},\cdots,u_{i-1}]$ denotes the tokenized texts before the response part, and $n$ represents the maximum length of output tokens.
Figure 5: Qualitative comparison with existing methods in CXR-to-report generation on the MIMIC-CXR (1st row) and OpenI datasets (2nd row). The texts in different colors show similar meanings. The keywords on the right are obtained from AdaMatch model.
Similarly, for report-to-CXR generation, we feed the LLM with the instruction $I^{v}$, the generated report $x^{T^{\prime}}$, and the image tokens of keypatches $E(k^{I})$ in the instruction-following format, and obtain the conditional probability of $k$-th token $w_{k}$, $P(w_{k})=\mathcal{LLM}(I^{v},x^{T^{\prime}},E(k^{I}))$. The CXR generation loss $L_{img}$ is defined as:
$L_{img}=\sum_{i=k}^{n}-\log P(w_{i}|w_{1},w_{2},\cdots,w_{i-1}),$ (11)
With $L_{text}$ and $L_{img}$, the LLM can implement the bidirectional CXR and report generation according to instructions.
Table 1: CXR-to-report generation performance in comparison to existing methods on MIMIC-CXR and OpenI datasets.
MIMIC-CXR OpenI
Methods BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L
R2Gen [7] (2020) 0.3553 0.2232 0.1523 0.1038 0.1412 0.2784 0.3992 0.2407 0.1518 0.0973 0.1390 0.3052
R2GenCMN [8] (2021) 0.3719 0.2332 0.1538 0.1053 0.1501 0.2827 0.4091 0.2493 0.1594 0.1045 0.1509 0.3181
Yang et al. [48] (2021) 0.3585 0.2266 0.1550 0.1021 0.1425 0.2788 0.3833 0.2409 0.1598 0.1078 0.1457 0.3293
XProNet [42] (2022) 0.3532 0.2212 0.1498 0.1052 0.1415 0.2811 0.4114 0.2502 0.1598 0.1045 0.1457 0.3240
ITHN [39] (2023) 0.3623 0.2128 0.1402 0.0992 0.1488 0.2622 0.2661 0.1516 0.0976 0.0663 0.1561 0.2617
M2KT [47] (2023) 0.3661 0.2192 0.1465 0.1044 0.1528 0.2673 0.2559 0.1381 0.0819 0.0523 0.1468 0.2439
AdaMatch-Cyclic 0.3793 0.2346 0.1540 0.1060 0.1625 0.2859 0.4161 0.3002 0.2073 0.1446 0.1621 0.3656
4 Experiments
4.1 Experiment Setting
Datasets. We conduct our experiments on two main publicly available chest X-ray datasets, i.e. MIMIC-CXR [20] and OpenI [11] datasets. MIMIC-CXR dataset is the largest public dataset for chest radiographs, including 473,057 chest X-ray images and 206,563 reports from 63,478 patients. We follow the official splits, i.e. 368,960 instances for training, 2,991 for validation, and 5,159 for testing. OpenI is a chest X-ray dataset with 3,684 report-image pairs, including 2,912 pairs for training and 772 pairs for testing. Different from existing methods [7, 8, 48, 47], we adopt both finding and impression sections as medical reports and remove the reports with less than 3 tokens for two datasets. There is no overlap of patients among different subsets. The CXR images with various sizes are all resized to $256\times 256$ pixels.
Implementation Details We adopt the PVT-medium [43] pre-trained with the disease classification task as the image encoder with AdaPatch in stage 2, 3 and 4. The AdaMatch is optimized with LAMB optimizer [51] and cosine learning rate scheduler [29]. The learning rate is 6e-3. The maximum training epoch is 15 and the per GPU batch size is 112. In AdaMatch-Cyclic, we adopt the VQ-GAN [13] models pre-trained on the MIMIC-CXR and OpenI datasets, respectively, and the dolly-v2-3b [9] model as pre-trained LLM. LLM has 5,1845 token types with 5,0821 token types for text tokens and 1,024 token types for image tokens. LLM is optimized by the AdamW [30] optimizer with a learning rate of 5e-6 and per GPU batch size of 24. The training epoch is 5. 8 Nvidia A100 40GB GPUs are used for all experiments. For more details, see supplementary materials.
Evaluation Metrics We evaluate the CXR-to-report generation performance with BLEU [33], METEOR [12], and ROUGE-L [27]. The FID [15] and NIQE [31] are utilized to evaluate report-to-CXR generation. The performance of CXR-to-report and report-to-CXR retrieval is assessed through the exact report in the top K retrieved reports for a given CXR image (R@K, K=$\left\{1,5,10\right\}$).
4.2 Comparison with State-of-the-Arts.
4.2.1 CXR-to-Report Generation
We verify the effectiveness of AdaMatch-Cyclic on the MIMIC-CXR and OpenI datasets for report generation. Since previous methods [7, 8, 48, 42, 39, 47] are mainly trained on the medical reports with the finding section, we reimplement previous methods on datasets with both finding and impression sections. In Table 1, our AdaMatch-Cyclic substantially surpasses the M2KT [47] by a significant margin with the BLEU-1 of 0.0132 in the MIMIC-CXR dataset. Moreover, we evaluate the OpenI dataset with the model trained on the MIMIC-CXR dataset. In terms of the zero-shot report generation, our AdaMatch-Cyclic achieves the best performance with the BLEU-1, METEOR, ROUGE-L of 0.4161, 0.1621, and 0.3656. To further compare with GPT-4V[49], we randomly select 50 cases from the test set of the MIMIC-CXR dataset. In Table 2, our AdaMatch-Cyclic significantly outperforms the GPT-4V by 0.1479 and 0.0892 of BLEU-1 and BLEU-4, indicating our superiority to existing methods.
Table 2: CXR-to-report generation performance in comparison to GPT-4V on 50 selected cases of the MIMIC-CXR dataset.
Methods BLEU-1 BLEU-2 BLEU-3 BLEU-4
R2Gen[7](2020) 0.3513 0.2174 0.1447 0.1025
R2GenCMN[8](2021) 0.3587 0.2200 0.1436 0.0969
Yang et al. [48](2021) 0.3596 0.2218 0.1481 0.1026
XProNet[42](2022) 0.3356 0.2071 0.1374 0.0941
ITHN[39](2023) 0.3301 0.1839 0.1121 0.0723
M2KT[47](2023) 0.3626 0.2123 0.1391 0.0957
GPT-4V[49](2023) 0.2275 0.0878 0.0378 0.0166
AdaMatch-Cyclic 0.3754 0.2303 0.1520 0.1058
Table 3: Report-to-CXR generation performance in comparison to existing methods on MIMIC-CXR and OpenI datasets.
MIMIC-CXR OpenI
[…]
Stable diffusion [37] 9.2334 3.7894 8.2946 6.3496
Chambon et al. [3] 8.2758 3.8871 5.8557 4.6534
RoentGen [2] 9.5411 3.8834 6.5675 4.9085
UniXGen [22] 6.7212 3.7125 11.9890 4.6610
LLM-CXR [23] 2.1788 3.5969 1.6597 3.8206
AdaMatch-Cyclic 1.0916 3.3931 1.5938 3.3096
To qualitatively evaluate report generation performance, we show two cases of generated reports in Fig. 5. In comparison to M2KT [47] and R2GenCMN [8], our AdaMatch-Cyclic can obtain the related keywords (e.g. scarring and atelectasis) and correctly describes the ‘osseous abnormality’ and ‘clips in the right upper quadrant of the abdomen’ in the first case, and point out ‘mild granulomatous’ in the second case, suggesting that our method can generate medical reports with higher quality than previous methods.
4.2.2 Report-to-CXR Generation
To quantitatively assess the performance, we compare our method with the text-to-image generation method [37] and some report-to-CXR generation methods [3, 2, 22, 23]. As listed in Table 3, AdaMatch-Cyclic achieves the best performance with 1.0916 and 1.5938 of FID scores on the MIMIC-CXR and OpenI datasets, respectively, indicating the superiority of our method over previous methods. To compare the high-level feature distribution of CXR images generated by different methods, we randomly select 1,000 cases from the test set, and apply the t-SNE visualization to the real and synthetic CXR images on the MIMIC-CXR dataset. As depicted in Fig. 6, synthetic CXR images generated by existing methods and real ones differ significantly from each other, while CXR images synthesized by AdaMatch-Cyclic almost overlap with the real CXR images. It implies that AdaMatch-Cyclic is more effective in translating medical reports to realistic CXR images compared to existing methods.
Figure 6: The t-SNE visualization of the real and synthetic CXR images on the MIMIC-CXR dataset.
Figure 7: The generated chest X-ray images of the MIMIC-CXR (1st row) and OpenI (2nd row) datasets with highlighted regions.
In Fig. 7, we show the generated CXR images of the MIMIC-CXR and OpenI datasets. In the first example, our AdaMatch-Cyclic can synthesize the ‘left and right pleural effusions’ as highlighted in the CXR images, while other methods [3, 2, 37] cannot generate the CXR images with ‘pleural effusions’. It suggests that our method can generate more realistic CXR images than existing methods according to the input report.
4.2.3 CXR-Report Retrieval
To prove the superiority of AdaMatch, we evaluate the CXR-to-report and report-to-CXR retrieval performance on the MIMIC-CXR datasets in comparison with existing methods. As listed in Table 4, our AdaMatch achieves the best retrieval performance, surpassing the LIMITR [10] by a large margin with R@1 of 11.77% and 13.48% for CXR-to-report and report-to-CXR retrieval, respectively, implying that our method is more effective in extracting distinct semantic features for CXR images and medical reports and is able to align their features correctly.
Table 4: CXR-report retrieval performance (%) in comparison to existing methods on the MIMIC-CXR dataset.
CXR-to-Report Report-to-CXR
Methods R@1 R@5 R@10 R@1 R@5 R@10
Fang et al. [14] (2015) 18.60 43.10 56.10 18.13 43.20 55.97
Chauhan et al. [4] (2020) 5.37 19.43 30.73 5.40 20.23 30.23
ConVIRT [53] (2020) 30.10 53.90 63.80 29.20 54.70 64.40
GLoRIA [16] (2021) 30.30 57.50 66.50 24.00 51.80 62.80
JoImTeR-Net [18] (2021) 18.93 46.20 58.67 19.07 45.27 58.50
MGCA [41] (2022) 25.80 51.90 62.10 27.90 51.20 61.60
LIMITR [10] (2023) 39.70 63.20 71.70 37.70 62.10 71.30
Motor [26] (2023) 10.96 31.93 42.90 12.00 33.10 44.32
AdaMatch 51.47 86.19 94.77 51.18 86.46 94.60
Table 5: Ablation study on AdaMatch w/o AdaPatch.
CXR-to-Report Report-to-CXR
AdaPatch R@1 R@5 R@10 R@1 R@5 R@10
48.77 83.89 92.94 48.72 83.95 92.90
51.47 86.19 94.77 51.18 86.46 94.60
4.3 Ablation Study
Effectiveness of AdaPatch. To prove the effectiveness of AdaPatch, we compare the CXR-report retrieval performance of AdaMatch model with or without AdaPatch. In the AdaMatch model without AdaPatch, we use PVT-medium [43] as the image encoder. In Table 7, the AdaPatch module remarkably improves the R@1 by 2.70% and 2.46% for CXR-to-report and report-to-CXR retrieval, respectively, indicating the effectiveness of our AdaPatch.
Image Encoder with Different Stages. In AdaMatch, the image encoder consists of four stages, and the stage 2,3 and 4 include an AdaPatch module, respectively. To investigate the effectiveness of each stage for CXR-report alignment, we use patch embeddings from different stages to compute the CXR-report similarity for CXR-report retrieval task. In Table 6, the image encoder with three stages achieves the best retrieval performance with the R@1 of 51.47%, implying that the features of the 3rd stage are more beneficial for patch-word matching.
Table 6: CXR-report retrieval performance of AdaMatch with different stages.
CXR-to-Report Report-to-CXR
Stage R@1 R@5 R@10 R@1 R@5 R@10
2 42.28 78.75 89.85 42.81 78.96 89.67
3 51.47 86.19 94.77 51.18 86.46 94.60
4 49.68 83.04 92.67 49.39 83.09 92.84
Visualization of Adaptive Patches. To visualize adaptive patches in AdaPatch, we display several examples from the MIMIC-CXR datasets. In Fig. 8, each example includes the text, its text tokens, and the corresponding CXR image with the marked adaptive patches. The adaptive patches are marked by the bounding boxes with different colors. Each bounding box includes an arrow, which shows the shift of center from the fixed patch to the adaptive patch. In the first example, adaptive patches cover the pacemaker and its wire. Meanwhile, adaptive patches of the second example show the correct position of pleural effusions in the CXR image. These indicate that AdaPatch can correctly localize the region relevant to input text tokens.
Figure 8: Visualization of texts and the corresponding adaptive patches. The boxes and arrows in different colors show adaptive patches and their shift of center from fixed patches.
5 Conclusion
We propose an AdaMatch for fine-grained image-text alignment, which presents the first work to adaptively associate image patches with words and improve the explainability of cyclic CXR-report generation. It includes the AdaPatch to acquire adaptive patches for abnormal regions, and performs patch-word alignment between adaptive patches with textual tokens. We implement cyclic CXR-report generation by using AdaMatch to provide explanations for the generation process. In addition, the fine-grained cyclic generation process supplies a natural explainability for the alignment between medical images and reports. Extensive experiments on two publicly available CXR datasets prove the effectiveness of our method and its superiority over previous methods.
Supplementary Material
In this supplementary material, we provide additional details to augment the content of the main paper. In Section 6, we conduct the ablation studies on the proposed AdaMatch-Cyclic. In Section 7, we demonstrate the visual results for CXR-to-report and report-to-CXR generation, and also visualize the adaptive patches and texts. In Section 8, we show implementation details of the proposed method.
6 Ablation Studies on AdaMatch-Cyclic
In the proposed AdaMatch-Cyclic, we leverage the AdaMatch to obtain keywords for CXR-to-report generation and keypatches for the report-to-CXR generation. To analyze the effectiveness of keywords and keypatches for the generation process, we ablate them in the AdaMatch-Cyclic on the MIMIC-CXR dataset to evaluate the CXR-report generation performance.
Effectiveness of keywords. As listed in Table 7, we ablate both keywords and keypatches of AdaMatch-Cyclic to create a baseline model. When we employ the keywords in the baseline model, the CXR-to-report generation performance improves significantly by about 0.03 in BLEU-4, indicating the effectiveness of the keywords obtained from AdaMatch. To further analyze the influence of the number of keywords $N_{w}$, we train the proposed AdaMatch-Cyclic with different $N_{w}=\left\{10,15,20\right\}$. In Table 8, AdaMatch-Cyclic achieves the best report generation performance with the ROUGE-L of 0.2859, when the $N_{w}$ is set to 10.
Effectiveness of keypatches. In Table 7, when we further apply keypatches to the baseline model with keywords, the CXR generation performance boosts remarkably by about 0.5 in FID score, implying the effectiveness of keypatches provided by AdaMatch. To investigate the influence of the number of keypatches $N_{p}$, we analyze the CXR generation performance when the number of keypatches ranges from 5 to 15. As shown in Table 9, AdaMatch-Cyclic achieves the best CXR generation performance with the FID score of 1.0916, when the number of keypatches is 5.
7 Visual Results
Table 7: The effectiveness of keywords and keypatches.
Keywords Keypatches BLEU-4 METEOR ROUGE-L FID
0.0778 0.1295 0.2402 1.5378
0.1059 0.1649 0.2813 1.5132
0.1060 0.1625 0.2859 1.0916
Table 8: The analysis on the different number of keywords ($N_{w}$).
$N_{w}$ BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L
10 0.3793 0.2346 0.1540 0.1060 0.1625 0.2859
15 0.3849 0.2265 0.1424 0.0937 0.1502 0.2668
20 0.3769 0.2281 0.1478 0.1001 0.1519 0.2762
Table 9: The analysis on the different number of keypatches($N_{p}$).
$N_{p}$ 5 10 15
FID 1.0916 1.6544 1.6720
Figure 9: Qualitative comparison with existing methods in CXR-to-report generation on the MIMIC-CXR (1st row) and OpenI (2nd row) datasets. The texts in different colors show similar meanings. The keywords on the right are obtained from AdaMatch model.
Figure 10: Visualization of the real and synthetic CXR images of the MIMIC-CXR (1st row) and the OpenI (2nd row) datasets.
Figure 11: Visualization of texts and the corresponding adaptive patches. The boxes and arrows in different colors show adaptive patches and their shift of center from fixed patches.
To prove the effectiveness of AdaMatch-Cyclic for CXR-to-report and report-to-CXR generation tasks, we demonstrate some visual results for both tasks. Moreover, to further visualize the explanation provided by AdaMatch, i.e. the adaptive patches and the corresponding texts.
CXR-to-Report Generation. As shown in Fig. 9, we compare CXR-to-report generation performance with existing methods on the MIMIC-CXR and OpenI datasets. In the first row, the proposed AdaMatch-Cyclic can capture the ‘Feeding tube tip’ and ‘mild bibasilar atelectasis’, while M2KT [47] cannot observe such device and lung abnormality. The second case of the OpenI dataset shows that ‘cardiac silhouette is mildly enlarged’ and ‘degenerative changes of the thoracic spine’ can be discovered by our method. These imply that our AdaMatch-Cyclic can generate a more comprehensive and complete medical report with the guidance of generated keywords in comparison to current methods.
Report-to-CXR Generation. Fig. 10 visualizes the real and synthetic CXR images of the MIMIC-CXR and OpenI datasets in comparison with existing methods. As depicted in the first example, the CXR image generated by AdaMatch-Cyclic shows ‘patchy opacity in the right lung’, while the CXR images generated by other methods do not include this pattern. In the second example, our AdaMatch-Cyclic can generate the CXR image with ‘low lung volumes’. These indicate the superiority of our AdaMatch-Cyclic over existing methods in report-to-CXR generation.
Adaptive Patches and Texts. In Fig. 11, we visualize adaptive patches in AdaPatch, the textual words, and the textual tokens for different cases from the MIMIC-CXR datasets. We highlight the adaptive patches with bounding boxes in different colors. Each bounding box has an arrow inside to show the shift of center from the fixed patch to the adaptive patch. In the first example, adaptive patches cover the pulmonary edema. Meanwhile, adaptive patches of the second example show the correct position of PICC device in the left lung of the CXR image. These suggest that AdaMatch-Cyclic can show the correspondence between the adaptive patches and textual words to provide the correct explanation for the CXR-report generation.
8 Implementation Details
In the AdaMatch-Cyclic model, we first train the AdaMatch model and then use the frozen AdaMatch to train LLM. The AdaMatch model consists of an image encoder and a text encoder. We utilize the PVT-medium [43] as the image encoder that includes AdaPatch module in stage 2, 3, and 4. The image encoder is pre-trained on the MIMIC-CXR dataset with the disease classification task. We adopt the pre-trained BioClinicalBERT [1] as text encoder. The image and text encoders are followed by two convolutional layers with batch normalization and the ReLU activation function, respectively, to reduce the feature dimension to 256. The patch size $s$ is set to 4 for stage 1, and 2 for stage 2, 3, and 4. The number of sampled feature points $m$ is 3. We optimize the AdaMatch using LAMB optimizer [51] ($\beta_{1}=0.9,\beta_{2}=0.999,\epsilon={10}^{-4}$). The cosine learning rate scheduler [29] with the base learning rate ($lr\_b$) of $6\times 10^{-3}$ is adopted to linearly warm up to the peak learning rate (i.e. $lr\_p=lr\_b\times\sqrt{\frac{bs\_all}{512}}$) during the first quarter of the total training epochs, where $bs\_all$ denotes the effective total batch size. The total training epoch is 15 and the per GPU batch size is 112. In AdaMatch-Cyclic, we use the pre-trained VQ-GAN [13] model to encode CXR images into image tokens and decode image tokens into CXR images. We adopt the dolly-v2-3b [9] model as the pre-trained LLM. The LLM has 5,1845 token types with the first 5,0821 token types for text tokens and the rest 1,024 token types for image tokens. We add 5,0821 to each image token value encoded by VQ-GAN. To decode the image tokens into images, we subtract 5,0821 from the image tokens generated by LLM and feed image tokens to VQ-GAN decoder to obtain the generated CXR images. We train the LLM with the AdamW [30] optimizer. The learning rate is initialized as $5\times{10}^{-6}$ and the total training epoch is 5. The per GPU batch size is set to 24. The hyper-parameters $\kappa_{0},\kappa_{1},\kappa_{2},\kappa_{3},\kappa_{4}$ are set as 200, 1, 1000, 20, and 5, respectively. In CXR-to-report generation, we extract keywords for each adaptive patch and use the keywords with the top 10 patch-word similarities as hints of LLM. In addition, we use 5 keypatches to guide the LLM in report-to-CXR generation. All the experiments are conducted on 8 Nvidia A100 40GB GPUs. The source code will be released.

### v5

@@
-Fine-Grained Image-Text Alignment in Medical Imaging Enables Cyclic Image-Report Generation
+Fine-Grained Image-Text Alignment in Medical Imaging Enables Explainable Cyclic Image-Report Generation
 Abstract
 Fine-grained vision-language models (VLM) have been widely used for inter-modality local alignment between the predefined fixed patches and textual words.
-However, in the medical analysis, lesions exhibit varying sizes and positions, and the fixed patches may cause incomplete representations of lesions.
-Moreover, these methods provide explainability by using heatmaps to show the general potential image areas associated with texts rather than specific regions, making their explanations not explicit and specific enough.
+However, in medical analysis, lesions exhibit varying sizes and positions, and using fixed patches may cause incomplete representations of lesions.
+Moreover, these methods provide explainability by using heatmaps to show the general image areas potentially associated with texts rather than specific regions, making their explanations not explicit and specific enough.
 To address these issues, we propose a novel Adaptive patch-word Matching (AdaMatch) model to correlate chest X-ray (CXR) image regions with words in medical reports and apply it to CXR-report generation to provide explainability for the generation process.
 AdaMatch exploits the fine-grained relation between adaptive patches and words to provide explanations of specific image regions with corresponding words.
-To capture the abnormal regions of varying sizes and positions, we introduce the Adaptive Patch extraction (AdaPatch) module to acquire the adaptive patches for these regions adaptively.
-In order to provide explicit explainability for CXR-report generation task, we propose an AdaMatch-based bidirectional large language model for Cyclic CXR-report generation (AdaMatch-Cyclic).
-It employs the AdaMatch to obtain the keywords for CXR images and ‘keypatches’ for medical reports as hints to guide CXR-report generation.
-Extensive experiments on two publicly available CXR datasets prove the effectiveness of our method and its superior performance to existing methods.
+To capture the abnormal regions of varying sizes and positions, we introduce an Adaptive Patch extraction (AdaPatch) module to acquire adaptive patches for these regions adaptively.
+Aiming to provide explicit explainability for the CXR-report generation task, we propose an AdaMatch-based bidirectional LLM for Cyclic CXR-report generation (AdaMatch-Cyclic).
+It employs AdaMatch to obtain the keywords for CXR images and ‘keypatches’ for medical reports as hints to guide CXR-report generation.
+Extensive experiments on two publicly available CXR datasets validate the effectiveness of our method and its superior performance over existing methods.
+Fine-Grained Image-Text Alignment in Medical Imaging
+Enables Explainable Cyclic Image-Report Generation
+Wenting Chen1  Linlin Shen3  Jingyang Lin4  Jiebo Luo4 Xiang Li5  Yixuan Yuan2thanks: Xiang Li and Yixuan Yuan are corresponding authors.
+1City University of Hong Kong   2The Chinese University of Hong Kong 3Shenzhen University   4University of Rochester 5Massachusetts General Hospital and Harvard Medical School 1wentichen7-c@my.cityu.edu.hk  2yxyuan@ee.cuhk.edu.hk  3llshen@szu.edu.cn 4{jluo@cs, jlin81@ur}.rochester.edu 5xli60@mgh.harvard.edu
 1 Introduction
-Inter-modality alignment, such as vision and language, has been an important task with growing interests in the field of computer vision, especially with the recent advancement in representation learning [34].
+Inter-modality alignment, such as vision and language, has been an important task with growing interests in the field of computer vision, especially with the recent advancement in representation learning Radford et al.
+(2021).
 Technologies like contrastive learning and self-supervised learning have dramatically improved state-of-the-art alignment performance.
-Recent vision-language models (VLMs) demonstrate two approaches: global contrastive alignment, which integrates images and texts at a global level [34, 19, 17, 44, 46], and local alignment, focusing on detailed connections between visual objects and textual words [6, 25, 24, 52, 21, 50], as illustrated in Fig.
+Recent vision-language models (VLMs) demonstrate two approaches: global contrastive alignment, which integrates images and texts at a global level Radford et al.
+(2021); Jia et al.
+(2021); Jang et al.
+(2023); Wang et al.
+(2023); Yang et al.
+(2022), and local alignment, focusing on detailed connections between visual objects and textual words Chen et al.
+(2020a); Li et al.
+(2020b, a); Zhan et al.
+(2021); Kim et al.
+(2021); Yao et al.
+(2021), as illustrated in Fig.
 1.
-Figure 1: Current vision-language models (VLM) achieve (a) global alignment and (b) local alignment by matching overall visual with textual features, and aligning patch with word features, respectively.
-(c) To exploit the relation between textual words and abnormal patches with different sizes, the proposed AdaMatch obtains adaptive patch features and aligns them with word features.
-Current VLMs with local alignment either adopt the pre-trained object detector to extract region-of-interest (ROI) features from images and match the corresponding object features with textual words [6, 25, 24, 52]; or align the visual token from each patch and the textual token into the same embedding space [21, 50, 18, 41].
+Figure 1: Current vision-language models (VLM) achieve (a) global alignment and (b) local alignment by matching overall visual with textual features, and aligning patches with word features, respectively.
+(c) To exploit the relation between textual words and abnormal patches with varied sizes, our AdaMatch obtains adaptive patch features and aligns them with word features.
+Current VLMs with local alignment either adopt the pre-trained object detector to extract region-of-interest (ROI) features from images and match the corresponding object features with textual words Chen et al.
+(2020a); Li et al.
+(2020b, a); Zhan et al.
+(2021), or align the visual token from each patch and the textual token into the same embedding space Kim et al.
+(2021); Yao et al.
+(2021); Ji et al.
+(2021); Wang et al.
+(2022a).
 The former highly relies on the quality of the object detector and its predefined classes, which is less generalizable to new domains.
 The latter family of methods learns the alignment in a more automatic and data-driven manner.
-However, most of these methods depend on pre-defined patch size and position (e.g., grids) across images.
-In most challenging cases, such as the analysis of medical image analysis, lesions can exhibit a wide range of shapes, sizes, and positions.
-A fixed definition of image patches can lead to incomplete or ambiguous representations of the key imaging abnormalities.
-Hence, it is highly needed to adaptively exploit the fine-grained relationship between image embeddings derived from a more flexible patching scheme and textual embeddings.
-Another challenge in the current local medical VLMs lies in their explainability: it is generally difficult to delineate the image-text relationship learned by the model, especially for the current medical VLMs.
-Current solutions to provide such explanations in medical VLMs leverage the attention maps from the intermediate layer to visualize the location of the abnormalities [32, 16, 45].
-Other methods  [40, 5, 28] utilize network gradients such as Grad-CAM [38] to generate the heatmaps according to lesion types based on ground-truth reports.
-However, both maps can only show the general potential areas associated with the corresponding text data rather than clearly pinpointing a specific region.
+However, most of these methods depend on a pre-defined patch size and positions (e.g., grids) across images.
+In the most challenging cases, such as the analysis of medical image, lesions can exhibit a wide range of shapes, sizes, and positions.
+A fixed partition of image patches can lead to incomplete or ambiguous representations of the key imaging abnormalities.
+Therefore, it is highly desirable to adaptively exploit the fine-grained relationship between image embeddings derived from a more flexible patching scheme and textual embeddings.
+Another challenge in the current VLMs lies in their explainability: it is generally difficult to delineate the image-text relationship learned by the model, especially for the current medical VLMs.
+Current solutions to provide such explanations in medical VLMs leverage the attention maps from the intermediate layer to visualize the location of the abnormalities Moon et al.
+(2022); Huang et al.
+(2021); Yan and Pei (2022).
+Other methods  Wan et al.
+(2023); Chen et al.
+(2023); Liu et al.
+(2023) utilize network gradients such as Grad-CAM Selvaraju et al.
+(2017) to generate the heatmaps according to lesion types based on ground-truth reports.
+However, both maps can only show the general areas potentially associated with the corresponding text data rather than pinpointing a specific region.
 In addition, gradient-based methods need ground-truth reports, prohibiting them from functioning correctly beyond training data.
-It is thus highly necessary to develop a mechanism that could provide explicit and specific explanations of the input image or text during inference time.
+It is thus highly necessary to develop a mechanism that could provide explicit and specific explanations of input image or text during inference time.
 Figure 2: The overview of the proposed methods.
@@
 Our premise is that such a cyclic generation task would serve as the best use case and evaluation criterion for the desired fine-grained alignment.
-Also, the fine-grained cyclic generation between CXR and report will provide a natural explainability for how the model aligns the two modalities: for any given text token, we can visualize its matching imaging manifestation; and for any image region within a CXR image, we can tell the type of lesion or anatomical region it belongs to.
+Also, fine-grained cyclic generation between CXR and report will provide natural explainability for how the model aligns two modalities: for any given text token, we can visualize its matching imaging manifestation; and for any image region within a CXR image, we can tell the type of lesion or anatomical region it belongs to.
 To implement the cyclic CXR-report generation, we propose an AdaMatch-based bidirectional model (AdaMatch-Cyclic).
-AdaMatch-Cyclic employs AdaMatch to identify the keywords for CXR images and ‘keypatches’ for medical reports to guide the generation tasks.
+AdaMatch-Cyclic employs AdaMatch to identify the keywords for CXR images and the ‘keypatches’ for medical reports to guide the generation tasks.
 Since the potential keywords for CXR images cover a wide range and ground-truth reports cannot be used during inference, we predefine a textual codebook with the most common entities from medical reports as prior knowledge during fine-grained alignment.
-With the textual codebook, AdaMatch aligns it with the adaptive patches to obtain matched keywords as guidance for report generation.
-Next, a VQ-GAN model encodes the CXR image into image tokens, and a Large Language Model (LLM) takes image tokens, the matched keywords, and the instructions as input to generate the medical reports.
+With the textual codebook, AdaMatch aligns it with the adaptive patches to obtain matched keywords to facilitate report generation.
+Next, a VQ-GAN model encodes the CXR image into image tokens, and a Large Language Model (LLM) takes image tokens, the matched keywords, and the instructions as input to generate medical reports.
 Similarly, we also build a visual codebook with the most commonly seen patches as ‘keypatches’, and use AdaMatch to obtain the matched keypatches from given text reports as hints for CXR generation.
-With medical reports, matched keypatches, and instructions, LLM outputs the image tokens, and then the VQ-GAN model decodes them into the generated CXR image.
+Utilizing medical reports, matched keypatches, and instructions, LLM generates image tokens, subsequently decoded by the VQ-GAN model to produce the resulting CXR image.
 Our contributions are summarized as follows:
-  • To exploit the fine-grained relation between the CXR image patches with words of medical reports, we propose an Adaptive patch-word Matching (AdaMatch) model to obtain adaptive patches for abnormal regions and perform alignment between them and texts in medical reports.
-  • We devise an AdaMatch-based bidirectional LLM for Cyclic CXR-report generation (AdaMatch-Cyclic) to facilitate the generation between CXR and reports.
+  • To exploit the fine-grained relation between CXR image patches and words of medical reports, we propose an Adaptive patch-word Matching (AdaMatch) model to obtain adaptive patches for abnormal regions and perform alignment between them and texts in medical reports.
+  • We devise an AdaMatch-based bidirectional LLM for Cyclic CXR-report generation (AdaMatch-Cyclic) to facilitate the bi-directional generation between CXR and reports.
 Moreover, we build the textual and visual codebook to utilize AdaMatch to extract useful keywords and keypatches for the report and CXR generation, respectively.
-  • Experiments on two publicly available chest X-ray datasets demonstrate the effectiveness of our method and its superior performance over state-of-art methods.
+  • Experiments on two publicly available chest X-ray datasets demonstrate the effectiveness of our method and its superior performance over the state-of-art methods.
 2 Related Works
 2.1 Fine-Grained Vision-Language Models
-Recently, several fine-grained vision-language models (VLM) [6, 25, 24, 52, 21, 50, 18, 41] achieve the local alignment by exploiting the fine-grained relation between visual objects and textual words.
-Some methods [6, 25, 24, 52] employ the pre-trained object detector to obtain region-of-interest (ROI) features from images and align the object features with textual features, and others [21, 50, 41, 18] aim to locally align the fixed patches with the textual words.
-For instance, Li et al.
-[24] utilizes Faster-RCNN [36] to detect visual objects and match ROI embeddings with textual embeddings.
-Moreover, FILIP [50] achieves cross-modal interaction by maximizing the similarity between the visual tokens of fixed patches and textual tokens.
-The former requires a precise object detector, and the latter primarily learns the relation between fixed patches inside the predefined grid and words.
-Nevertheless, when lung lesions differ in size, shape, and position, these methods may split them into separate patches, leading to incomplete semantic information of patches.
+Recently, several fine-grained vision-language models (VLM) achieve the local alignment by exploiting the fine-grained relation between visual objects and textual words.
+Some methods Chen et al.
+(2020a); Li et al.
+(2020b, a); Zhan et al.
+(2021) employ the pre-trained object detector to obtain object features from images and align them with textual features, and others Kim et al.
+(2021); Yao et al.
+(2021); Wang et al.
+(2022a); Ji et al.
+(2021); Xu et al.
+(2023); Jiang et al.
+(2023) aim to align the fixed patches with the textual words locally.
+The former relies on a precise object detector, and the latter focuses on the relation between fixed patches inside the predefined grid and words.
+However, varying lung lesion characteristics may cause these methods to divide them into separate patches, resulting in incomplete semantic information.
 Thus, we devise an Adaptive patch-word Matching (AdaMatch) model to adaptively exploit the fine-grained relation between flexible patches and textual words.
 2.2 CXR-Report Generation
-Medical VLMs are widely used in downstream tasks for chest radiographs, including CXR-to-report generation [7, 8, 48, 42, 39, 47] and report-to-CXR generation [37, 3, 2, 22, 23] tasks.
-CXR-to-report generation has been investigated to reduce the burden of radiologists and promote clinical automation.
-For instance, Chen et al.
-[8] introduces a cross-modal memory network with shared memory to align images with texts to promote report generation performance.
-In report-to-CXR generation task, previous methods generate high-quality and annotated CXR images from medical reports, to provide more training data and avoid privacy laws and considerations.
-Some diffusion-based methods [2, 22] take the impression section of reports as input, ignoring the finding section that includes detailed descriptions.
-To consider more details in reports, several transformer-based methods [22, 23] take both finding and impression sections as input to synthesize CXR.
+Medical VLMs are widely used in downstream tasks for chest radiographs, including CXR-to-report generation Chen et al.
+(2020b, 2021a); Yang et al.
+(2021); Wang et al.
+(2022b); Voutharoja et al.
+(2023); Yang et al.
+(2023); Shi et al.
+(2023); Huang et al.
+(2023) and report-to-CXR generation Rombach et al.
+(2022); Chambon et al.
+(2022b, a); Lee et al.
+(2023a, b); Han et al.
+(2024); Shentu and Al Moubayed (2024); Hou et al.
+(2023); Hashmi et al.
+(2024); Chen et al.
+(2024) tasks.
+For CXR-to-report generation, Chen et al., 2021a introduces a cross-modal memory network with shared memory to align images with texts to promote report generation performance.
+In the report-to-CXR generation task, prior techniques create annotated CXR images from medical reports to augment training data and address privacy concerns, which are categorized into diffusion-based and transformer-based methods.
 In this paper, we use the cyclic generation (i.e.
@@
 In Fig.
-2, we propose an adaptive patch-word matching (AdaMatch) model to associate the CXR image regions with words in medical reports, and apply it to CXR-report generation to provide explainability.
+2, we propose an adaptive patch-word matching (AdaMatch) model to associate CXR image regions with words in medical reports, and apply it to CXR-report generation to enhance explainability.
 Given an input CXR image $x^{I}$, the image encoder with several Adaptive Patch extraction modules (AdaPatch) predicts the location and scale of multiple adaptive patches and processes the adaptive patch embeddings $z^{I}$.
-With $z^{I}$ and text embeddings $z^{T}$ extracted by text encoder for the medical report $x^{T}$, we compute the similarities $s^{I}$, $s^{T}$ between adaptive patches and text tokens to calculate the contrastive loss $L_{I2T},L_{T2I}$ to optimize AdaMatch model.
+With $z^{I}$ and text embeddings $z^{T}$ extracted by text encoder for the medical report $x^{T}$, we compute the similarities $s^{I}$, $s^{T}$ between adaptive patches and text tokens to calculate the contrastive loss $L_{I2T},L_{T2I}$ to optimize AdaMatch.
 For CXR-to-report generation, we utilize the frozen AdaMatch model to match a predefined textual codebook with the input CXR image $x^{I}$ to obtain matched keywords $k^{T}$ and feed a LLM with $k^{T}$, the instruction, and image tokens encoded by VQ-GAN from $x^{I}$ to generate a medical report $x^{T^{\prime}}$.
 Then, the AdaMatch model is used to match a predefined visual codebook with the generated report $x^{T^{\prime}}$ to acquire the matched keypatches $k^{I}$.
-With instruction, $x^{T^{\prime}}$ and $k^{I}$, LLM outputs image tokens of the generated CXR image $x^{I^{\prime}}$, which is optimized by $L_{text}$ and $L_{img}$.
+With instruction, $x^{T^{\prime}}$ and $k^{I}$, LLM outputs image tokens of the generated CXR image $x^{I^{\prime}}$, which is optimized through $L_{text}$ and $L_{img}$.
 3.1 Adaptive Patch-Word Matching (AdaMatch)
-Current VLMs [6, 25, 24, 52, 21, 50] achieve fine-grained alignment between visual objects and textual words, but they may split lung lesions into separate fixed patches due to various shapes, sizes, and positions of lung lesions.
-Thus, we propose an Adaptive patch-word Matching (AdaMatch) to automatically locate the important regions, extract adaptive patches for these regions, and align them with corresponding textual words.
-As shown in Fig.
-2 (a), AdaMatch includes an image encoder to extract adaptive patch features and a text encoder to encode textual features.
+Current VLMs Chen et al.
+(2020a); Yao et al.
+(2021) achieve fine-grained alignment between visual objects and textual words, but they may split lung lesions into separate fixed patches due to various sizes of lung lesions.
+Thus, we propose an Adaptive patch-word Matching (AdaMatch) to locate the important regions, extract adaptive patches for these regions, and align them with corresponding textual words, as shown in Fig.
+2 (a).
 Image Encoder.
-To obtain the adaptive patches, the image encoder consists of four stages with feature maps of decreasing scales.
-Specifically, in the first stage, we first adopt the patch embedding module to split the CXR image $x^{I}\in\mathbb{R}^{H\times W\times C}$ into $N$ patches with fixed size $s\times s$ and project them to patch embeddings $z^{(i)}(1\leq i\leq N)$ through a fully connected layer $g$, $z^{(i)}=g([a^{(i,1)};\dots;a^{(i,s\times s)}])$, where $a^{(i,j)}$ indicates the image features for the pixel located at $q^{(i,j)}$ and []delimited-[][$\cdot$][ ⋅ ] denotes the concatenation operation.
-Afterward, we pass the patch embeddings $z^{(i)}$ with the position embeddings $e_{pos}$ into a transformer encoder [43] $R$ to obtain the outputs, $z^{(i)}=R(z^{(i)}+e_{pos})$, where $R$ includes an attention layer and a feed-forward layer.
-In order to locate the potential lung lesions, we devise an Adaptive Patch Extraction module (AdaPatch) for the rest of stages.
-In AdaPatch, a fully connected layer is fed with patch embeddings $z^{(i)}$, and then the outputs are passed through four separate fully connected layers $f_{1},f_{2},f_{3},f_{4}$ to predict the offset $(\delta_{x}^{(i)},\delta_{y}^{(i)})$ and patch size $(s_{w}^{(i)},s_{h}^{(i)})$ for the adaptive patches, where the offset indicates the shift to the center of original fixed patches $(c^{(i)}_{x},c^{(i)}_{y})$,
-$\delta_{x}^{(i)}=\small{Tanh}(f_{1}(z^{(i)})),\delta_{y}^{(i)}=\small{Tanh}(f_{2}(z^{(i)})),$ (1)
-$s_{w}^{(i)}=\small{ReLU}(\small{Tanh}(f_{3}(z^{(i)})),s_{h}^{(i)}=\small{ReLU}(\small{Tanh}(f_{4}(z^{(i)})).$ (2)
+To obtain the adaptive patches, the image encoder comprises four stages with feature maps of decreasing scales.
+Specifically, in the first stage, we first adopt the patch embedding module to split the CXR image $x^{I}\in\mathbb{R}^{H\times W\times C}$ into $N$ patches with fixed size $s\times s$ and project them to patch embeddings $z^{(i)}(1\leq i\leq N)$ through a fully connected (FC) layer $g$, $z^{(i)}=g([a^{(i,1)};\dots;a^{(i,s\times s)}])$, where $a^{(i,j)}$ indicates the image features for the pixel located at the $q^{(i,j)}$.
+$q^{(i,j)}$ indicates the coordinates for the image pixel.
+[]delimited-[][$\cdot$][ ⋅ ] represents the concatenation operation among the features.
+Afterward, we pass the patch embeddings $z^{(i)}$ with the position embeddings $e_{pos}$ into a transformer encoder $R$ to obtain the outputs, $z^{(i)}=R(z^{(i)}+e_{pos})$.
+To locate the potential lung lesions, we devise an Adaptive Patch Extraction module (AdaPatch) for the rest of stages.
+In AdaPatch, patch embeddings $z^{(i)}$ are fed into an FC layer, followed by four separate FC layers $f_{1},f_{2},f_{3},f_{4}$ to predict offsets $(\delta_{x}^{(i)},\delta_{y}^{(i)})$ and patch sizes $(s_{w}^{(i)},s_{h}^{(i)})$ for adaptive patches, with offsets indicating shifts to the center $(c^{(i)}_{x},c^{(i)}_{y})$ of fixed patches,
+$\delta_{x}^{(i)}={Tanh}(f_{1}(z^{(i)})),\delta_{y}^{(i)}={Tanh}(f_{2}(z^{(i)})),$ (1)
+$s_{w}^{(i)}={ReLU}({Tanh}(f_{3}(z^{(i)})),s_{h}^{(i)}={ReLU}({Tanh}(f_{4}(z^{(i)})).$ (2)
 With the offset and patch size, we compute the position of left-top $(a^{(i)}_{x},a^{(i)}_{y})$ and right-bottom corners $(b^{(i)}_{x},b^{(i)}_{y})$ for each adaptive patch,
@@
 Since the coordinates may be fractional, bilinear interpolation is used to obtain the sampled feature points $\left\{\hat{p}^{(j)}\right\}_{1\leq j\leq m\times m}$.
-The embeddings of all sampled points are flattened and fed into a fully connected layer $f_{5}$ to obtain patch embeddings, $z^{(i)}=f_{5}([\hat{p}^{(i,1)};\dots;\hat{p}^{(i,m\times m)}])$.
+The embeddings of all sampled points are flattened and fed into an FC layer $f_{5}$ to obtain patch embeddings, $z^{(i)}=f_{5}([\hat{p}^{(i,1)};\dots;\hat{p}^{(i,m\times m)}])$.
 Finally, we pass $z^{(i)}$ with position embeddings $e_{pos}$ into a transformer encoder $R$.
-The final adaptive patch embeddings $z^{I}\in\mathbb{R}^{N\times d}$ are the ensemble of $z^{(i)}(1\leq i\leq N)$, $z^{I}=E^{I}(x^{I})=\left\{z^{I}_{1},\cdots,z^{I}_{N}\right\},$ where $E^{I}$, $x^{I}$, and $d$ denote the image encoder, the input CXR image, and the dimension of adaptive patch embeddings, respectively.
+The final adaptive patch embeddings $z^{I}\in\mathbb{R}^{N\times d}$ are the ensemble of $z^{(i)}(1\leq i\leq N)$, $z^{I}=E^{I}(x^{I})=\left\{z^{I}_{1},\cdots,z^{I}_{N}\right\},$ where $E^{I}$, $x^{I}$, and $d$ denote the image encoder, the CXR image, and the dimension of $z^{I}$, respectively.
 Text Encoder.
-We adopt the BioClinicalBERT [1] as the text encoder $E^{T}$ to encode the medical report $x^{T}$ into text embeddings $z^{T}$.
-To make $z^{T}$ with the same dimension as $z^{I}$, we feed a fully connected layer with $z^{T}$ to reduce its dimension, $z^{T}=E^{T}(x^{T})=\left\{z^{T}_{1},\cdots,z^{T}_{K}\right\}(z^{T}\in\mathbb{R}^{K\times d})$, where $K$ denotes the number of text tokens.
+We adopt a pre-trained text encoder $E^{T}$ to encode the medical report $x^{T}$ into text embeddings $z^{T}$.
+To make $z^{T}$ with the same dimension as $z^{I}$, we feed an FC layer with $z^{T}$ to reduce its dimension, $z^{T}=E^{T}(x^{T})=\left\{z^{T}_{1},\cdots,z^{T}_{K}\right\}(z^{T}\in\mathbb{R}^{K\times d})$, where $K$ is the number of text tokens.
 Patch-Word Alignment.
-In order to exploit the relation between the adaptive patches and textual tokens, we perform fine-grained contrastive representation learning to achieve patch-word alignment.
+To exploit the relation between the adaptive patches and textual tokens, we perform fine-grained contrastive representation learning to achieve patch-word alignment.
 Concretely, for the $i$-th CXR image $x^{I}_{i}$ and $j$-th medical report $x^{T}_{j}$, we first compute the similarities between all the adaptive patch embeddings $z^{I}_{n}(1\leq n\leq N)$ and all the text embeddings $z^{T}_{k}(1\leq k\leq K)$, and use the largest similarity $\max_{(1\leq k\leq K)}{{(z^{I}_{n})}^{\top}z^{T}_{k}}$ as the patch-word maximum similarity for $n$-th adaptive patch embedding.
-Then, the patch-word maximum similarities for all the adaptive patch embeddings are averaged as the similarity $s^{I}_{i,j}$ of the $i$-th CXR image to the $j$-th medical report,
-$s^{I}_{i,j}=\frac{1}{N}\displaystyle\sum_{n=1}^{N}(z^{I}_{n})^{\top}z^{T}_{m_{n}^{I}},$ (5)
-where $m_{n}^{I}=\arg\max_{(1\leq k\leq K)}{{(z^{I}_{n})}^{\top}z^{T}_{k}}$. Similarly, the similarity of the $j$-th medical report to the $i$-th CXR image is defined as,
-$s^{T}_{i,j}=\frac{1}{K}\displaystyle\sum_{k=1}^{K}(z^{I}_{m_{k}^{T}})^{\top}z^{T}_{k},$ (6)
-where $m_{k}^{T}=\arg\max_{(1\leq n\leq N)}{{(z^{I}_{n})}^{\top}z^{T}_{k}}$. We exclude the padded textual tokens when computing the similarity.
+Then, the patch-word maximum similarities for all the adaptive patch embeddings are averaged as the similarity $s^{I}_{i,j}$ of the $i$-th CXR image to the $j$-th medical report, $s^{I}_{i,j}=\frac{1}{N}\displaystyle\sum_{n=1}^{N}(z^{I}_{n})^{\top}z^{T}_{m_{n}^{I}}$, where $m_{n}^{I}=\arg\max_{(1\leq k\leq K)}{{(z^{I}_{n})}^{\top}z^{T}_{k}}$.
+Similarly, the similarity of the $j$-th medical report to the $i$-th CXR image is defined as, $s^{T}_{i,j}=\frac{1}{K}\displaystyle\sum_{k=1}^{K}(z^{I}_{m_{k}^{T}})^{\top}z^{T}_{k}$, where $m_{k}^{T}=\arg\max_{(1\leq n\leq N)}{{(z^{I}_{n})}^{\top}z^{T}_{k}}$.
+We exclude the padded textual tokens when computing the similarity.
 With the cross-modal similarities $s^{I}_{i,j}$ and $s^{T}_{i,j}$ for the $i$-th CXR image and $j$-th medical report, we compute the adaptive patch-word contrastive loss $L_{i}^{I}$ for $i$-th CXR image $x_{i}^{I}$,
 $L_{i}^{I2T}(x_{i}^{I},\left\{x_{j}^{T}\right\}_{j=1}^{b})=-\frac{1}{b}\log\frac{exp(s^{I}_{i,i}/\tau)}{\textstyle\sum_{j}{exp(s_{i,j}^{I}/\tau)}},$ (5)
-where $b$ denotes the batch size, $\tau$ represents temperature hyperparameter and $\left\{x_{i}^{I},x_{i}^{T}\right\}$ indicates the positive CXR-report pair.
+where $b$, $\tau$, and $\left\{x_{i}^{I},x_{i}^{T}\right\}$ represent the batch size, temperature hyperparameter and the positive CXR-report pair, respectively.
 Similarly, the adaptive word-patch contrastive loss $L_{i}^{T}$ for $i$-th medical report $x_{i}^{T}$ is formulated as,
 $L_{i}^{T2I}(x_{i}^{T},\left\{x_{j}^{I}\right\}_{j=1}^{b})=-\frac{1}{b}\log\frac{exp(s^{T}_{i,i}/\tau)}{\textstyle\sum_{j}{exp(s_{j,i}^{T}/\tau)}}.$ (6)
-The final contrastive loss for a mini-batch is calculated by,
-$L=\frac{1}{2}\sum_{i=1}^{b}{(L_{i}^{I2T}+L_{i}^{T2I})}.$ (9)
-With the final contrastive loss, the image and text encoders are optimized to automatically locate the important patches with varied sizes and exploit the fine-grained relation between the patches and words.
+The final contrastive loss for a mini-batch is calculated by, $L=\frac{1}{2}\sum_{i=1}^{b}{(L_{i}^{I2T}+L_{i}^{T2I})}$.
+With $L$, AdaMatch learns to locate important patches with varied sizes and exploits the fine-grained relation.
 3.2 AdaMatch-based LLM for Cyclic CXR-Report Generation (AdaMatch-Cyclic)
@@
 2 (b).
-Figure 3: Instruction data example for CXR-to-report generation.
 3.2.1 CXR-to-Report Generation
-In CXR-to-report generation, we employ the AdaMatch to match a predefined textual codebook with the CXR image to obtain the matched keywords as hints to guide LLM to generate reports.
+In CXR-to-report generation, we use AdaMatch to match a predefined textual codebook with the CXR image, providing keywords to guide LLM.
 Building textual codebook.
-Specifically, to establish a textual codebook, we use a pre-trained BioEN [35] to extract the related entities from the medical reports in the training set, where the related entities are divided into four entity groups, i.e., biological structure, detailed description, disease disorder, and sign symptom.
+Specifically, we first use a pre-trained BioEN Raza et al.
+(2022) to extract the related entities from medical reports in the training set, where the related entities are divided into four entity groups, i.e., biological structure, detailed description, disease disorder, and sign symptom.
 Next, we compute the frequency of each entity and pick top $\kappa_{0}$ entities for each entity group as keywords in the textual codebook.
 Keywords Extraction.
-With the textual codebook, we employ the frozen AdaMatch model to match the keywords from the textual codebook with the adaptive patches of CXR images, and obtain a patch-word similarity matrix $s^{P}\in\mathbb{R}^{M\times N}$ between the keyword tokens and adaptive patch embeddings, where $M$ and $N$ denote the number of keyword tokens and adaptive patches.
-To extract the most matched keywords for each adaptive patch, we rank the patch-word similarity along the dimension of keyword tokens, obtain the top $\kappa_{1}$ patch-word similarities for each adaptive patch, and extract the corresponding keywords $k^{T}$.
+With the textual codebook, we employ the frozen AdaMatch model to match keywords from the textual codebook with the adaptive patches of CXR images, and obtain a patch-word similarity matrix $s^{P}\in\mathbb{R}^{M\times N}$ between the keyword tokens and adaptive patch embeddings, where $M$ and $N$ denote the number of keyword tokens and adaptive patches.
+To extract the most matched keywords for each adaptive patch, we rank the patch-word similarity along the dimension of keyword tokens, obtain the top $\kappa_{1}$ patch-word similarities for each adaptive patch, and extract corresponding keywords $k^{T}$.
 The matched keywords can explain potential lesions in each adaptive patch, to assist LLM in generating medical reports.
-Figure 4: Instruction data example for report-to-CXR generation.
+Figure 3: The example of instruction data for CXR-to-report generation.
 Instruction tuning LLM.
-After keywords extraction, we use a frozen VQ-GAN [13] encoder $E$ to encode the input CXR image $x^{I}$ to the quantized image latent vectors as image tokens $E(x^{I})$.
-For CXR-to-report generation, the dolly-v2-3b [9] model is adopted as the bidirectional LLM, which is pre-trained on the instruction-following dataset.
-When fine-tuning LLM, we convert CXR-to-report generation dataset in instruction-following format.
-An example of instruction data for CXR-to-report generation is shown in Fig.
-3, including the instruction, input, keywords, and response parts, where input is the image tokens of the input CXR image, keywords are extracted by AdaMatch, and response is the ground-truth medical report.
-During training, the response part is hidden and the LLM learns to generate it in autoregressive manner.
-By adopting AdaMatch-Cyclic for CXR-to-report generation, we can locate the potential lesion in CXR images, to provide more explainability for the generation process.
+After keywords extraction, we use a frozen VQ-GAN Esser et al.
+(2021) encoder $E$ to encode the input CXR image $x^{I}$ to the quantized image latent vectors as image tokens $E(x^{I})$.
+For CXR-to-report generation, we adopt the dolly-v2-3b Conover et al.
+(2023) model as the pre-trained LLM, and convert CXR-to-report generation dataset in instruction-following format.
+An example of instruction data for CXR-to-report generation is depicted in Fig.
+3, with the instruction, input (the image tokens of the input CXR image), keywords (extracted by AdaMatch), and response parts (the ground-truth medical report).
+During training, the LLM learns to generate the hidden response part in an autoregressive manner.
+By adopting AdaMatch-Cyclic, we can locate potential lesions and interpret them with keywords to enhance explainability.
 3.2.2 Report-to-CXR Generation
-In report-to-CXR generation, we leverage the AdaMatch to match a predefined visual codebook with the generated reports $x^{T^{\prime}}$ to acquire matched keypatches as guidance for LLM to synthesize CXR images, where ‘keypatches’ are important image patches related to generated reports.
+In report-to-CXR generation, AdaMatch matches a predefined visual codebook with generated reports $x^{T^{\prime}}$, providing keypatches as guidance for the LLM to synthesize CXR images, where keypatches are important image patches related to the reports.
 Building visual codebook.
 Concretely, we first construct a visual codebook with the most common adaptive patches as keypatches.
-To collect the most common adaptive patches for CXR images in the training set, the AdaMatch is utilized to match the adaptive patches of CXR images with textual tokens of medical reports to obtain top $\kappa_{2}$ CXR-report pairs with the highest report-to-CXR similarities $s^{T}$.
+To collect the most common adaptive patches for CXR images in the training set, AdaMatch matches the adaptive patches of CXR images with textual tokens of medical reports to obtain top $\kappa_{2}$ CXR-report pairs with the highest report-to-CXR similarities $s^{T}$.
 For each CXR-report pair, we compute the word-patch maximum similarity $\max_{(1\leq n\leq N)}{{(z^{I}_{n})}^{\top}z^{T}_{k}}$ for each textual token, rank the word-patch maximum similarities, and extract the adaptive patches for top $\kappa_{3}$ similarities as keypatches in the visual codebook.
-Each keypatch includes its adaptive patches and the corresponding features.
+Each keypatch includes its adaptive patch and the corresponding features.
 Keypatches Extraction.
-With the visual codebook, the frozen AdaMatch matches the features of keypatches in the visual codebook with textual tokens of the generated report to acquire the word-patch similarity matrix $s^{W}\in\mathbb{R}^{(\kappa_{2}\times\kappa_{3})\times K}$, where $K$ denotes the number of textual tokens for the generated report.
+With the visual codebook, the frozen AdaMatch matches the features of keypatches in the visual codebook with textual tokens of the generated report to acquire the word-patch similarity matrix $s^{W}\in\mathbb{R}^{(\kappa_{2}\times\kappa_{3})\times K}$, where $K$ is the number of textual tokens.
 To obtain keypatches related to the generated report, we rank the word-patch similarity along the dimension of keypatches, obtain the top $\kappa_{4}$ word-patch similarity for each textual token, and extract the features of corresponding keypatches $k^{I}$.
+Table 1: Comparison of CXR-to-report generation performance on the MIMIC-CXR and the OpenI datasets.
+MIMIC-CXR OpenI
+Methods B-1 B-2 B-3 B-4 M R-L B-1 B-2 B-3 B-4 M R-L
+R2Gen 0.3553 0.2232 0.1523 0.1038 0.1412 0.2784 0.3992 0.2407 0.1518 0.0973 0.1390 0.3052
+R2GenCMN 0.3719 0.2332 0.1538 0.1053 0.1501 0.2827 0.4091 0.2493 0.1594 0.1045 0.1509 0.3181
+Joint-TriNet 0.3585 0.2266 0.1550 0.1021 0.1425 0.2788 0.3833 0.2409 0.1598 0.1078 0.1457 0.3293
+XProNet 0.3532 0.2212 0.1498 0.1052 0.1415 0.2811 0.4114 0.2502 0.1598 0.1045 0.1457 0.3240
+ITHN 0.3623 0.2128 0.1402 0.0992 0.1488 0.2622 0.2661 0.1516 0.0976 0.0663 0.1561 0.2617
+M2KT 0.3661 0.2192 0.1465 0.1044 0.1528 0.2673 0.2559 0.1381 0.0819 0.0523 0.1468 0.2439
+AdaMatch-Cyclic 0.3793 0.2346 0.1540 0.1060 0.1625 0.2859 0.4161 0.3002 0.2073 0.1446 0.1621 0.3656
 Instruction tuning LLM.
-After keypatches extraction, we employ a frozen VQ-GAN encoder to convert the matched keypatches $k^{I}$ into image tokens $E(k^{I})$, and feed the pre-trained LLM with the instruction, the generated report, and image tokens of keypatches in the instruction-following format, as shown in Fig.
+After keypatches extraction, we use a frozen VQ-GAN encoder to convert the matched keypatches $k^{I}$ into image tokens $E(k^{I})$, and feed the LLM with the instruction, generated report, and image tokens of keypatches in the instruction-following format, as shown in Fig.
 4.
-Then, LLM predicts the image tokens and the VQ-GAN decoder decodes the image tokens into the corresponding generated CXR image $x^{I^{\prime}}$.
+Then, LLM predicts image tokens, decoded by the VQ-GAN into the generated CXR image $x^{I^{\prime}}$.
+Figure 4: The example of instruction data for report-to-CXR generation.
 AdaMatch-Cyclic allows us to interpret generated reports with matched keypatches, thereby providing explainability to the generation procedure.
 3.2.3 Overall Objective
-To optimize AdaMatch-Cyclic, we apply the standard language modeling objective for both CXR-to-report and report-to-CXR generation.
-For CXR-to-report generation, LLM is fed with the instruction $I^{t}$, image tokens of input CXR image $E(x^{I})$, and matched keywords $k^{T}$ in instruction-following format, and LLM aims to generate the ground-truth medical report of response part autoregressively.
-We compute the $P(u_{k})$ conditional probability of the $k$-th token, $P(u_{k})=\mathcal{LLM}(I^{t},E(x^{I}),k^{T})$, where $k$ is the token index after the response key (### Response:).
-The report generation loss $L_{text}$ for the response area is calculated by,
-$L_{text}=\sum_{i=k}^{n}-\log P(u_{i}|u_{1},u_{2},\cdots,u_{i-1}),$ (10)
-where $[u_{1},u_{2},\cdots,u_{i-1}]$ denotes the tokenized texts before the response part, and $n$ represents the maximum length of output tokens.
-Figure 5: Qualitative comparison with existing methods in CXR-to-report generation on the MIMIC-CXR (1st row) and OpenI datasets (2nd row).
-The texts in different colors show similar meanings.
-The keywords on the right are obtained from AdaMatch model.
-Similarly, for report-to-CXR generation, we feed the LLM with the instruction $I^{v}$, the generated report $x^{T^{\prime}}$, and the image tokens of keypatches $E(k^{I})$ in the instruction-following format, and obtain the conditional probability of $k$-th token $w_{k}$, $P(w_{k})=\mathcal{LLM}(I^{v},x^{T^{\prime}},E(k^{I}))$.
-The CXR generation loss $L_{img}$ is defined as:
-$L_{img}=\sum_{i=k}^{n}-\log P(w_{i}|w_{1},w_{2},\cdots,w_{i-1}),$ (11)
-With $L_{text}$ and $L_{img}$, the LLM can implement the bidirectional CXR and report generation according to instructions.
-Table 1: CXR-to-report generation performance in comparison to existing methods on MIMIC-CXR and OpenI datasets.
-MIMIC-CXR OpenI
-Methods BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L
-R2Gen [7] (2020) 0.3553 0.2232 0.1523 0.1038 0.1412 0.2784 0.3992 0.2407 0.1518 0.0973 0.1390 0.3052
-R2GenCMN [8] (2021) 0.3719 0.2332 0.1538 0.1053 0.1501 0.2827 0.4091 0.2493 0.1594 0.1045 0.1509 0.3181
-Yang et al. [48] (2021) 0.3585 0.2266 0.1550 0.1021 0.1425 0.2788 0.3833 0.2409 0.1598 0.1078 0.1457 0.3293
-XProNet [42] (2022) 0.3532 0.2212 0.1498 0.1052 0.1415 0.2811 0.4114 0.2502 0.1598 0.1045 0.1457 0.3240
-ITHN [39] (2023) 0.3623 0.2128 0.1402 0.0992 0.1488 0.2622 0.2661 0.1516 0.0976 0.0663 0.1561 0.2617
-M2KT [47] (2023) 0.3661 0.2192 0.1465 0.1044 0.1528 0.2673 0.2559 0.1381 0.0819 0.0523 0.1468 0.2439
-AdaMatch-Cyclic 0.3793 0.2346 0.1540 0.1060 0.1625 0.2859 0.4161 0.3002 0.2073 0.1446 0.1621 0.3656
+To optimize AdaMatch-Cyclic, we use the standard language modeling objective for both CXR-to-report and report-to-CXR generation.
+For CXR-to-report generation, LLM receives the instruction $I^{t}$, image tokens of input CXR image $E(x^{I})$, and matched keywords $k^{T}$, aiming to generate the ground-truth medical report of response part autoregressively.
+We compute the conditional probability of the $k$-th token, $P(u_{k})=\mathcal{LLM}(I^{t},E(x^{I}),k^{T})$, where $k$ is the token index after the response key (### Response:).
+The report generation loss for the response area is calculated by, $L_{text}=\sum_{i=k}^{n}-\log P(u_{i}|u_{1},u_{2},\cdots,u_{i-1})$, where $[u_{1},u_{2},\cdots,u_{i-1}]$ denotes the tokenized texts before the response part, and $n$ represents the maximum length of output tokens.
+Similarly, for report-to-CXR generation, we feed the LLM with the instruction $I^{v}$, generated report $x^{T^{\prime}}$, and image tokens of keypatches $E(k^{I})$, and obtain the conditional probability of $k$-th token $w_{k}$, $P(w_{k})=\mathcal{LLM}(I^{v},x^{T^{\prime}},E(k^{I}))$.
+The CXR generation loss is defined as: $L_{img}=\sum_{i=k}^{n}-\log P(w_{i}|w_{1},w_{2},\cdots,w_{i-1})$.
+With $L_{text}$ and $L_{img}$, LLM can implement bidirectional CXR and report generation according to instructions.
+Figure 5: Qualitative comparison of CXR-to-report generation on the MIMIC-CXR (1st row) and the OpenI (2nd row) datasets, highlighting similar meanings in colored text.
+The keywords are obtained from AdaMatch.
 4 Experiments
@@
 Datasets.
-We conduct our experiments on two main publicly available chest X-ray datasets, i.e.
-MIMIC-CXR [20] and OpenI [11] datasets.
-MIMIC-CXR dataset is the largest public dataset for chest radiographs, including 473,057 chest X-ray images and 206,563 reports from 63,478 patients.
-We follow the official splits, i.e.
-368,960 instances for training, 2,991 for validation, and 5,159 for testing.
-OpenI is a chest X-ray dataset with 3,684 report-image pairs, including 2,912 pairs for training and 772 pairs for testing.
-Different from existing methods [7, 8, 48, 47], we adopt both finding and impression sections as medical reports and remove the reports with less than 3 tokens for two datasets.
-There is no overlap of patients among different subsets.
-The CXR images with various sizes are all resized to $256\times 256$ pixels.
-Implementation Details We adopt the PVT-medium [43] pre-trained with the disease classification task as the image encoder with AdaPatch in stage 2, 3 and 4.
-The AdaMatch is optimized with LAMB optimizer [51] and cosine learning rate scheduler [29].
-The learning rate is 6e-3.
-The maximum training epoch is 15 and the per GPU batch size is 112.
-In AdaMatch-Cyclic, we adopt the VQ-GAN [13] models pre-trained on the MIMIC-CXR and OpenI datasets, respectively, and the dolly-v2-3b [9] model as pre-trained LLM.
-LLM has 5,1845 token types with 5,0821 token types for text tokens and 1,024 token types for image tokens.
-LLM is optimized by the AdamW [30] optimizer with a learning rate of 5e-6 and per GPU batch size of 24.
-The training epoch is 5.
-8 Nvidia A100 40GB GPUs are used for all experiments.
-For more details, see supplementary materials.
-Evaluation Metrics We evaluate the CXR-to-report generation performance with BLEU [33], METEOR [12], and ROUGE-L [27].
-The FID [15] and NIQE [31] are utilized to evaluate report-to-CXR generation.
-The performance of CXR-to-report and report-to-CXR retrieval is assessed through the exact report in the top K retrieved reports for a given CXR image (R@K, K=$\left\{1,5,10\right\}$).
+We experiment on two main publicly available chest X-ray datasets, i.e.
+MIMIC-CXR Johnson et al.
+(2019) and OpenI Demner-Fushman et al.
+(2016) datasets.
+MIMIC-CXR dataset comprises 473,057 images and 206,563 reports from 63,478 patients.
+We use the official splits, i.e.
+368,960 for training, 2,991 for validation, and 5,159 for testing.
+OpenI contains 3,684 report-image pairs with 2,912 for training and 772 for testing.
+Unlike prior methods, we utilize both finding and impression sections as medical reports.
+Table 2: Performance of CXR-to-report generation compared to GPT-4V on 50 selected MIMIC-CXR cases.
+Methods B-1 B-2 B-3 B-4
+R2Gen 0.3513 0.2174 0.1447 0.1025
+R2GenCMN 0.3587 0.2200 0.1436 0.0969
+Joint-TriNet 0.3596 0.2218 0.1481 0.1026
+XProNet 0.3356 0.2071 0.1374 0.0941
+ITHN 0.3301 0.1839 0.1121 0.0723
+M2KT 0.3626 0.2123 0.1391 0.0957
+GPT-4V 0.2275 0.0878 0.0378 0.0166
+AdaMatch-Cyclic 0.3754 0.2303 0.1520 0.1058
+Table 3: Comparison of report-to-CXR generation performance on the MIMIC-CXR and the OpenI datasets.
+MIMIC-CXR OpenI
+[…]
+Stable diffusion 9.2334 3.7894 8.2946 6.3496
+Adapting-Med 8.2758 3.8871 5.8557 4.6534
+RoentGen 9.5411 3.8834 6.5675 4.9085
+UniXGen 6.7212 3.7125 11.9890 4.6610
+LLM-CXR 2.1788 3.5969 1.6597 3.8206
+AdaMatch-Cyclic 1.0916 3.3931 1.5938 3.3096
+Implementation Details.
+In the AdaMatch-Cyclic model, we first train the AdaMatch model and then use the frozen AdaMatch to train LLM.
+We adopt VQ-GAN Esser et al.
+(2021) models pre-trained on the MIMIC-CXR and OpenI datasets, respectively, and the dolly-v2-3b Conover et al.
+(2023) model as pre-trained LLM.
+The source code will be released.
+Please see appendix C for more details.
+Evaluation Metrics.
+We assess CXR-to-report generation using BLEU (B), METEOR (M), and ROUGE-L (R-L) Chen et al.
+(2020b), and report-to-CXR generation using FID Heusel et al.
+(2017) and NIQE Mittal et al.
+(2012).
+The retrieval performance between CXR and report is assessed through the exact report in the top K retrieved reports for a given CXR image (R@K, K=$\left\{1,5,10\right\}$).
 4.2 Comparison with State-of-the-Arts.
 4.2.1 CXR-to-Report Generation
-We verify the effectiveness of AdaMatch-Cyclic on the MIMIC-CXR and OpenI datasets for report generation.
-Since previous methods [7, 8, 48, 42, 39, 47] are mainly trained on the medical reports with the finding section, we reimplement previous methods on datasets with both finding and impression sections.
-In Table 1, our AdaMatch-Cyclic substantially surpasses the M2KT [47] by a significant margin with the BLEU-1 of 0.0132 in the MIMIC-CXR dataset.
-Moreover, we evaluate the OpenI dataset with the model trained on the MIMIC-CXR dataset.
-In terms of the zero-shot report generation, our AdaMatch-Cyclic achieves the best performance with the BLEU-1, METEOR, ROUGE-L of 0.4161, 0.1621, and 0.3656.
-To further compare with GPT-4V[49], we randomly select 50 cases from the test set of the MIMIC-CXR dataset.
-In Table 2, our AdaMatch-Cyclic significantly outperforms the GPT-4V by 0.1479 and 0.0892 of BLEU-1 and BLEU-4, indicating our superiority to existing methods.
-Table 2: CXR-to-report generation performance in comparison to GPT-4V on 50 selected cases of the MIMIC-CXR dataset.
-Methods BLEU-1 BLEU-2 BLEU-3 BLEU-4
-R2Gen[7](2020) 0.3513 0.2174 0.1447 0.1025
-R2GenCMN[8](2021) 0.3587 0.2200 0.1436 0.0969
-Yang et al. [48](2021) 0.3596 0.2218 0.1481 0.1026
-XProNet[42](2022) 0.3356 0.2071 0.1374 0.0941
-ITHN[39](2023) 0.3301 0.1839 0.1121 0.0723
-M2KT[47](2023) 0.3626 0.2123 0.1391 0.0957
-GPT-4V[49](2023) 0.2275 0.0878 0.0378 0.0166
-AdaMatch-Cyclic 0.3754 0.2303 0.1520 0.1058
-Table 3: Report-to-CXR generation performance in comparison to existing methods on MIMIC-CXR and OpenI datasets.
-MIMIC-CXR OpenI
-[…]
-Stable diffusion [37] 9.2334 3.7894 8.2946 6.3496
-Chambon et al. [3] 8.2758 3.8871 5.8557 4.6534
-RoentGen [2] 9.5411 3.8834 6.5675 4.9085
-UniXGen [22] 6.7212 3.7125 11.9890 4.6610
-LLM-CXR [23] 2.1788 3.5969 1.6597 3.8206
-AdaMatch-Cyclic 1.0916 3.3931 1.5938 3.3096
-To qualitatively evaluate report generation performance, we show two cases of generated reports in Fig.
-5.
-In comparison to M2KT [47] and R2GenCMN [8], our AdaMatch-Cyclic can obtain the related keywords (e.g.
-scarring and atelectasis) and correctly describes the ‘osseous abnormality’ and ‘clips in the right upper quadrant of the abdomen’ in the first case, and point out ‘mild granulomatous’ in the second case, suggesting that our method can generate medical reports with higher quality than previous methods.
-4.2.2 Report-to-CXR Generation
-To quantitatively assess the performance, we compare our method with the text-to-image generation method [37] and some report-to-CXR generation methods [3, 2, 22, 23].
-As listed in Table 3, AdaMatch-Cyclic achieves the best performance with 1.0916 and 1.5938 of FID scores on the MIMIC-CXR and OpenI datasets, respectively, indicating the superiority of our method over previous methods.
-To compare the high-level feature distribution of CXR images generated by different methods, we randomly select 1,000 cases from the test set, and apply the t-SNE visualization to the real and synthetic CXR images on the MIMIC-CXR dataset.
-As depicted in Fig.
-6, synthetic CXR images generated by existing methods and real ones differ significantly from each other, while CXR images synthesized by AdaMatch-Cyclic almost overlap with the real CXR images.
-It implies that AdaMatch-Cyclic is more effective in translating medical reports to realistic CXR images compared to existing methods.
+We compare AdaMatch-Cyclic with current CXR-to-report generation methods on the MIMIC-CXR and the OpenI datasets, including R2Gen Chen et al.
+(2020b), R2GenCMN Chen et al.
+(2021a), Joint-TriNet Yang et al.
+(2021), XProNet Wang et al.
+(2022b), ITHN Voutharoja et al.
+(2023), and M2KT Yang et al.
+(2023).
+Since these methods are mainly trained on the finding section, we reimplement them on both the finding and impression sections.
+In Table 1, AdaMatch-Cyclic achieves the best performance of 0.3793 in BLEU-1 on the MIMIC-CXR dataset, with superior generalization to OpenI dataset using the same model.
+Compared to GPT-4V, AdaMatch-Cyclic significantly outperforms in BLEU-1 by 0.1479, as listed in Table 2.
 Figure 6: The t-SNE visualization of the real and synthetic CXR images on the MIMIC-CXR dataset.
-Figure 7: The generated chest X-ray images of the MIMIC-CXR (1st row) and OpenI (2nd row) datasets with highlighted regions.
-In Fig.
-7, we show the generated CXR images of the MIMIC-CXR and OpenI datasets.
-In the first example, our AdaMatch-Cyclic can synthesize the ‘left and right pleural effusions’ as highlighted in the CXR images, while other methods [3, 2, 37] cannot generate the CXR images with ‘pleural effusions’.
-It suggests that our method can generate more realistic CXR images than existing methods according to the input report.
-4.2.3 CXR-Report Retrieval
-To prove the superiority of AdaMatch, we evaluate the CXR-to-report and report-to-CXR retrieval performance on the MIMIC-CXR datasets in comparison with existing methods.
-As listed in Table 4, our AdaMatch achieves the best retrieval performance, surpassing the LIMITR [10] by a large margin with R@1 of 11.77% and 13.48% for CXR-to-report and report-to-CXR retrieval, respectively, implying that our method is more effective in extracting distinct semantic features for CXR images and medical reports and is able to align their features correctly.
-Table 4: CXR-report retrieval performance (%) in comparison to existing methods on the MIMIC-CXR dataset.
+Figure 7: Generated CXR images of the MIMIC-CXR (1st row) and OpenI (2nd row) datasets with highlighted regions.
+Table 4: Comparison of CXR-report retrieval performance (%) on MIMIC-CXR dataset.
 CXR-to-Report Report-to-CXR
 Methods R@1 R@5 R@10 R@1 R@5 R@10
-Fang et al. [14] (2015) 18.60 43.10 56.10 18.13 43.20 55.97
-Chauhan et al. [4] (2020) 5.37 19.43 30.73 5.40 20.23 30.23
-ConVIRT [53] (2020) 30.10 53.90 63.80 29.20 54.70 64.40
-GLoRIA [16] (2021) 30.30 57.50 66.50 24.00 51.80 62.80
-JoImTeR-Net [18] (2021) 18.93 46.20 58.67 19.07 45.27 58.50
-MGCA [41] (2022) 25.80 51.90 62.10 27.90 51.20 61.60
-LIMITR [10] (2023) 39.70 63.20 71.70 37.70 62.10 71.30
-Motor [26] (2023) 10.96 31.93 42.90 12.00 33.10 44.32
+Fang et al. (2015) 18.60 43.10 56.10 18.13 43.20 55.97
+Chauhan et al. (2020) 5.37 19.43 30.73 5.40 20.23 30.23
+ConVIRT Zhang et al. (2022) 30.10 53.90 63.80 29.20 54.70 64.40
+GLoRIA Huang et al. (2021) 30.30 57.50 66.50 24.00 51.80 62.80
+JoImTeR-Net Ji et al. (2021) 18.93 46.20 58.67 19.07 45.27 58.50
+MGCA Wang et al. (2022a) 25.80 51.90 62.10 27.90 51.20 61.60
+LIMITR Dawidowicz et al. (2023) 39.70 63.20 71.70 37.70 62.10 71.30
+Motor Lin et al. (2023) 10.96 31.93 42.90 12.00 33.10 44.32
 AdaMatch 51.47 86.19 94.77 51.18 86.46 94.60
@@
 51.47 86.19 94.77 51.18 86.46 94.60
-4.3 Ablation Study
-Effectiveness of AdaPatch.
-To prove the effectiveness of AdaPatch, we compare the CXR-report retrieval performance of AdaMatch model with or without AdaPatch.
-In the AdaMatch model without AdaPatch, we use PVT-medium [43] as the image encoder.
-In Table 7, the AdaPatch module remarkably improves the R@1 by 2.70% and 2.46% for CXR-to-report and report-to-CXR retrieval, respectively, indicating the effectiveness of our AdaPatch.
-Image Encoder with Different Stages.
-In AdaMatch, the image encoder consists of four stages, and the stage 2,3 and 4 include an AdaPatch module, respectively.
-To investigate the effectiveness of each stage for CXR-report alignment, we use patch embeddings from different stages to compute the CXR-report similarity for CXR-report retrieval task.
-In Table 6, the image encoder with three stages achieves the best retrieval performance with the R@1 of 51.47%, implying that the features of the 3rd stage are more beneficial for patch-word matching.
+Table 6: Effectiveness of report-to-CXR and CXR-to-report generation tasks.
+[…]
+0.3542 0.1973 0.1190 0.0758 0.1329 0.2392 - -
+- - - - - - 1.7128 4.0391
+0.3793 0.2346 0.1540 0.1060 0.1625 0.2813 1.0916 3.3931
+Table 7: Comparison of AdaPatch and other adaptive vision models.
+CXR-to-Report Report-to-CXR
+Models R@1 R@5 R@10 R@1 R@5 R@10
+A-ViT 49.78 83.14 92.77 49.49 83.19 92.94
+AdaViT 50.14 84.52 93.20 50.54 84.10 93.83
+AdaPatch 51.47 86.19 94.77 51.18 86.46 94.60
 Table 8: CXR-report retrieval performance of AdaMatch with different stages.
@@
 4 49.68 83.04 92.67 49.39 83.09 92.84
+Figure 8: Visualization of texts and its adaptive patches. Boxes and arrows in different colors show adaptive patches and their center shifts from fixed patches.
+In Fig.
+5, we compare AdaMatch-Cyclic’s performance with M2KT and R2GenCMN.
+AdaMatch-Cyclic accurately captures relevant keywords like ‘scarring’, and describes ‘osseous abnormality’ and ‘clips in the right upper quadrant of the abdomen’ in the first case, showcasing superior report quality over previous methods.
+4.2.2 Report-to-CXR Generation
+We quantitatively compare our method with text-to-image generation method like Stable Diffusion Rombach et al.
+(2022) and report-to-CXR generation methods, such as Adapting-Med Chambon et al.
+(2022b), RoentGen Chambon et al.
+(2022a), UniXGen Lee et al.
+(2023a), and LLM-CXR Lee et al.
+(2023b).
+In Table 3, AdaMatch-Cyclic achieves the highest FID scores on both datasets, showing its superior effectiveness in generating CXR.
+To compare the high-level feature distribution of CXR images generated by different methods, we randomly select 1,000 cases from test set, and apply t-SNE visualization to the real and synthetic CXR images on the MIMIC-CXR dataset.
+In Fig.
+6, while existing methods’ synthetic CXR images differ from real ones, AdaMatch-Cyclic’s almost overlap with real ones, indicating its superiority in report-to-CXR generation.
+In Fig.
+7, we display generated CXR images from the MIMIC-CXR and OpenI datasets.
+In the first example, AdaMatch-Cyclic accurately synthesizes ‘left and right pleural effusions’, while other methods fail to generate such features.
+This suggests the superior ability of our method to produce realistic CXR images based on input reports.
+4.2.3 CXR-Report Retrieval
+To demonstrate the superiority of AdaMatch, we evaluate the CXR-to-report and report-to-CXR retrieval performance on the MIMIC-CXR dataset compared to existing methods.
+We utilize the AdaMatch to compute the similarities between CXR images and reports, rank the similarities, and obtain the retrieval results.
+As shown in Table 4, AdaMatch significantly outperforms LIMITR with R@1 of 11.77% and 13.48% for CXR-to-report and report-to-CXR retrieval.
+This indicates the superior ability of our method to extract distinct semantic features and align them accurately between CXR images and medical reports.
+In Table 5, to ablate AdaPatch, we remove the layers to predict offset and scale.
+We then aligned grid features from the last stage of the pyramid vision Transformer with textual tokens to compute the image-text similarity for retrieval tasks.
+4.3 Ablation Study
+Effectiveness of AdaPatch.
+We evaluate the effectiveness of AdaPatch by comparing the CXR-report retrieval performance of AdaMatch with and without AdaPatch.
+To ablate AdaPatch, we remove the layers to predict offset and scale, and compute the similarity between grid features of the last stage and textual tokens for retrieval tasks.
+In Table 5, AdaPatch significantly improves R@1 by approximately 3% for both retrieval directions, highlighting its effectiveness.
+To demonstrate the optimality of the AdaPatch design, we compare AdaPatch with other adaptive vision models, such as A-ViT Yin et al.
+(2022) and AdaViT Meng et al.
+(2022), in the CXR-report retrieval task.
+As shown in Table 7, AdaPatch exhibits superior retrieval performance compared with other methods.
+Effectiveness of Cyclic Generation.
+To validate the necessity of report-to-CXR and CXR-to-report generation tasks, we remove each task and evaluate their performance, as listed in Table 6.
+When ablating report-to-CXR generation task, the performance of CXR-to-report generation decreases substantially compared to our method, and vice versa.
+These suggest that the CXR-to-report and report-to-CXR generation can indeed benefit each other, and their interaction is crucial for the overall performance of the model.
+Image Encoder with Different Stages.
+In AdaMatch, the image encoder comprises several stages, with AdaPatch modules in the stages.
+To assess the effectiveness of the stage number, we compare the retrieval performance of AdaMatch with two, three, or four stages.
+Table 8 shows that the three-stage image encoder achieves the highest R@1 of 51.47%, indicating that the features from the 3rd stage are most beneficial for CXR-report retrieval.
 Visualization of Adaptive Patches.
-To visualize adaptive patches in AdaPatch, we display several examples from the MIMIC-CXR datasets.
-In Fig.
-8, each example includes the text, its text tokens, and the corresponding CXR image with the marked adaptive patches.
-The adaptive patches are marked by the bounding boxes with different colors.
-Each bounding box includes an arrow, which shows the shift of center from the fixed patch to the adaptive patch.
-In the first example, adaptive patches cover the pacemaker and its wire.
-Meanwhile, adaptive patches of the second example show the correct position of pleural effusions in the CXR image.
-These indicate that AdaPatch can correctly localize the region relevant to input text tokens.
-Figure 8: Visualization of texts and the corresponding adaptive patches. The boxes and arrows in different colors show adaptive patches and their shift of center from fixed patches.
+We visualize adaptive patches in AdaPatch of some examples from the MIMIC-CXR dataset in Fig.
+8.
+Each example includes text, its tokens, and the corresponding CXR image with marked adaptive patches.
+The patches are represented by colored bounding boxes with arrows indicating the center shift from the fixed patch.
+In the first example, adaptive patches cover the pacemaker and wire, and the second one can find pleural effusions, implying the capacity of AdaPatch to localize regions relevant to input text tokens accurately.
 5 Conclusion
-We propose an AdaMatch for fine-grained image-text alignment, which presents the first work to adaptively associate image patches with words and improve the explainability of cyclic CXR-report generation.
-It includes the AdaPatch to acquire adaptive patches for abnormal regions, and performs patch-word alignment between adaptive patches with textual tokens.
+We propose AdaMatch for fine-grained image-text alignment, which presents the first work to adaptively associate image patches with words and improve the explainability of cyclic CXR-report generation.
+It includes AdaPatch to acquire adaptive patches for abnormal regions and performs patch-word alignment between adaptive patches with textual tokens.
 We implement cyclic CXR-report generation by using AdaMatch to provide explanations for the generation process.
-In addition, the fine-grained cyclic generation process supplies a natural explainability for the alignment between medical images and reports.
-Extensive experiments on two publicly available CXR datasets prove the effectiveness of our method and its superiority over previous methods.
-Supplementary Material
-In this supplementary material, we provide additional details to augment the content of the main paper.
-In Section 6, we conduct the ablation studies on the proposed AdaMatch-Cyclic.
-In Section 7, we demonstrate the visual results for CXR-to-report and report-to-CXR generation, and also visualize the adaptive patches and texts.
-In Section 8, we show implementation details of the proposed method.
-6 Ablation Studies on AdaMatch-Cyclic
+In addition, the fine-grained cyclic generation process provides a natural explainability for the alignment between CXR images and reports.
+Extensive experiments on two CXR datasets show the effectiveness of our method and its superiority over previous methods.
+Limitation
+While our research has made significant strides in utilizing chest X-ray datasets, it is important to acknowledge certain limitations.
+Our experiments predominantly focus on chest X-ray datasets due to their availability of large-scale images paired with high-quality medical reports.
+However, these datasets primarily consist of patients from the ICU, potentially skewing our model towards severe disease domains.
+In our future endeavors, we intend to expand the scope of our methodology by applying the proposed AdaMatch-Cyclic to multi-domain scenarios, thereby mitigating this limitation and enhancing the versatility of our approach.
+Acknowledge
+This work was supported by the Hong Kong Research Grants Council (RGC) General Research Fund under Grant 14220622, Innovation and Technology Commission Innovation and Technology Fund ITS/229/22, and the National Natural Science Foundation of China under Grant 82261138629.
+Appendix
+Abstract.
+In this supplementary material, we provide additional information about the proposed method.
+Appendix A illustrates ablation studies on AdaMatch-Cyclic.
+Appendix B demonstrates visual results of CXR-to-report generation, report-to-CXR generation, and adaptive patches.
+Appendix C provides the implementation details of the proposed method.
+Appendix A Ablation Studies on AdaMatch-Cyclic
 In the proposed AdaMatch-Cyclic, we leverage the AdaMatch to obtain keywords for CXR-to-report generation and keypatches for the report-to-CXR generation.
@@
 In Table 10, AdaMatch-Cyclic achieves the best report generation performance with the ROUGE-L of 0.2859, when the $N_{w}$ is set to 10.
+Table 9: The effectiveness of keywords and keypatches. B-4, M, and R-L represent BLEU-4, METEOR, and ROUGE-L, respectively.
+Keywords Keypatches B-4 M R-L FID
+0.0778 0.1295 0.2402 1.5378
+0.1059 0.1649 0.2813 1.5132
+0.1060 0.1625 0.2859 1.0916
+Table 10: The analysis on the different number of keywords ($N_{w}$). B, M, and R-L represent BLEU, METEOR, and ROUGE-L, respectively.
+$N_{w}$ B-1 B-2 B-3 B-4 M R-L
+10 0.3793 0.2346 0.1540 0.1060 0.1625 0.2859
+15 0.3849 0.2265 0.1424 0.0937 0.1502 0.2668
+20 0.3769 0.2281 0.1478 0.1001 0.1519 0.2762
 Effectiveness of keypatches.
@@
 As shown in Table 11, AdaMatch-Cyclic achieves the best CXR generation performance with the FID score of 1.0916, when the number of keypatches is 5.
-7 Visual Results
-Table 7: The effectiveness of keywords and keypatches.
-Keywords Keypatches BLEU-4 METEOR ROUGE-L FID
-0.0778 0.1295 0.2402 1.5378
-0.1059 0.1649 0.2813 1.5132
-0.1060 0.1625 0.2859 1.0916
-Table 8: The analysis on the different number of keywords ($N_{w}$).
-$N_{w}$ BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L
-10 0.3793 0.2346 0.1540 0.1060 0.1625 0.2859
-15 0.3849 0.2265 0.1424 0.0937 0.1502 0.2668
-20 0.3769 0.2281 0.1478 0.1001 0.1519 0.2762
 Table 11: The analysis on the different number of keypatches($N_{p}$).
@@
 Figure 10: Visualization of the real and synthetic CXR images of the MIMIC-CXR (1st row) and the OpenI (2nd row) datasets.
-Figure 11: Visualization of texts and the corresponding adaptive patches. The boxes and arrows in different colors show adaptive patches and their shift of center from fixed patches.
+Figure 11: Visualization of texts and the corresponding adaptive patches. The boxes and arrows in different colors show adaptive patches and their center shifts from fixed patches.
+Appendix B Visual Results
 To prove the effectiveness of AdaMatch-Cyclic for CXR-to-report and report-to-CXR generation tasks, we demonstrate some visual results for both tasks.
@@
 9, we compare CXR-to-report generation performance with existing methods on the MIMIC-CXR and OpenI datasets.
-In the first row, the proposed AdaMatch-Cyclic can capture the ‘Feeding tube tip’ and ‘mild bibasilar atelectasis’, while M2KT [47] cannot observe such device and lung abnormality.
+In the first row, the proposed AdaMatch-Cyclic can capture the ‘Feeding tube tip’ and ‘mild bibasilar atelectasis’, while M2KT Yang et al.
+(2023) cannot observe such device and lung abnormality.
 The second case of the OpenI dataset shows that ‘cardiac silhouette is mildly enlarged’ and ‘degenerative changes of the thoracic spine’ can be discovered by our method.
@@
 These suggest that AdaMatch-Cyclic can show the correspondence between the adaptive patches and textual words to provide the correct explanation for the CXR-report generation.
-8 Implementation Details
+Appendix C Implementation Details
 In the AdaMatch-Cyclic model, we first train the AdaMatch model and then use the frozen AdaMatch to train LLM.
 The AdaMatch model consists of an image encoder and a text encoder.
-We utilize the PVT-medium [43] as the image encoder that includes AdaPatch module in stage 2, 3, and 4.
+We utilize the DPT-medium Chen et al.
+(2021b) as the image encoder that includes AdaPatch module in stage 2 and 3.
 The image encoder is pre-trained on the MIMIC-CXR dataset with the disease classification task.
-We adopt the pre-trained BioClinicalBERT [1] as text encoder.
+We adopt the pre-trained BioClinicalBERT Alsentzer et al.
+(2019) as text encoder.
 The image and text encoders are followed by two convolutional layers with batch normalization and the ReLU activation function, respectively, to reduce the feature dimension to 256.
@@
 The number of sampled feature points $m$ is 3.
-We optimize the AdaMatch using LAMB optimizer [51] ($\beta_{1}=0.9,\beta_{2}=0.999,\epsilon={10}^{-4}$).
-The cosine learning rate scheduler [29] with the base learning rate ($lr\_b$) of $6\times 10^{-3}$ is adopted to linearly warm up to the peak learning rate (i.e.
+We optimize the AdaMatch using LAMB optimizer ($\beta_{1}=0.9,\beta_{2}=0.999,\epsilon={10}^{-4}$).
+The cosine learning rate scheduler with the base learning rate ($lr\_b$) of $6\times 10^{-3}$ is adopted to linearly warm up to the peak learning rate (i.e.
 $lr\_p=lr\_b\times\sqrt{\frac{bs\_all}{512}}$) during the first quarter of the total training epochs, where $bs\_all$ denotes the effective total batch size.
 The total training epoch is 15 and the per GPU batch size is 112.
-In AdaMatch-Cyclic, we use the pre-trained VQ-GAN [13] model to encode CXR images into image tokens and decode image tokens into CXR images.
-We adopt the dolly-v2-3b [9] model as the pre-trained LLM.
+In AdaMatch-Cyclic, we use the pre-trained VQ-GAN Esser et al.
+(2021) model to encode CXR images into image tokens and decode image tokens into CXR images.
+We adopt the dolly-v2-3b Conover et al.
+(2023) model as the pre-trained LLM.
 The LLM has 5,1845 token types with the first 5,0821 token types for text tokens and the rest 1,024 token types for image tokens.
@@
 To decode the image tokens into images, we subtract 5,0821 from the image tokens generated by LLM and feed image tokens to VQ-GAN decoder to obtain the generated CXR images.
-We train the LLM with the AdamW [30] optimizer.
+We train the LLM with the AdamW Loshchilov and Hutter (2018) optimizer.
 The learning rate is initialized as $5\times{10}^{-6}$ and the total training epoch is 5.
 The per GPU batch size is set to 24.
-The hyper-parameters $\kappa_{0},\kappa_{1},\kappa_{2},\kappa_{3},\kappa_{4}$ are set as 200, 1, 1000, 20, and 5, respectively.
+The hyper-parameters $\kappa_{0},\kappa_{1},\kappa_{2},\kappa_{3},\kappa_{4}$ are set as 200, 10, 1000, 20, and 5, respectively.
 In CXR-to-report generation, we extract keywords for each adaptive patch and use the keywords with the top 10 patch-word similarities as hints of LLM.
@@
 All the experiments are conducted on 8 Nvidia A100 40GB GPUs.
-The source code will be released.
+There is no overlap of patients among different subsets.
+We discard the medical reports with fewer than 3 tokens from both datasets.
+All the CXR images with different sizes are resized to $256\times 256$ pixels.